### v1

# Instruction-Based Molecular Graph Generation with Unified Text-Graph Diffusion Model

Abstract
Recent advancements in computational chemistry have increasingly focused on synthesizing molecules based on textual instructions. Integrating graph generation with these instructions is complex, leading most current methods to use molecular sequences with pre-trained large language models. In response to this challenge, we propose a novel framework, named UTGDiff (Unified Text-Graph Diffusion Model), which utilizes language models for discrete graph diffusion to generate molecular graphs from instructions. UTGDiff features a unified text-graph transformer as the denoising network, derived from pre-trained language models and minimally modified to process graph data through attention bias. Our experimental results demonstrate that UTGDiff consistently outperforms sequence-based baselines in tasks involving instruction-based molecule generation and editing, achieving superior performance with fewer parameters given an equivalent level of pretraining corpus. Our code is availble at https://github.com/ran1812/UTGDiff.
Introduction
Molecules, as fundamental units of matter, possess structures and properties that determine the characteristics of substances. Research into novel molecules is pivotal for chemistry, biology, and drug discovery (Schneider and Fechner 2005; Hughes et al. 2011), as these fields depend on the discovery of molecules with specific properties. A significant challenge in drug discovery is designing molecules that exhibit desired features, maintain structural stability, and interact with specific targets (Hartenfeller and Schneider 2011). Traditionally, this process has been resource-intensive and time-consuming (Dickson and Gagnon 2009; Khanna 2012). However, recent advancements in deep learning have revolutionized drug design by offering cost-effective methods (Nag et al. 2022; Askr et al. 2023), drawing significant attention to the research of generating drug-like molecules (Bagal et al. 2021; Schneuing et al. 2022).
In this work, we aim to bridge the gap between molecular structures and natural language, designing a generative model that functions akin to a drug design expert (illustrated in Figure 1). By receiving specific instructions, this model can precisely generate candidate molecules with the indicated characteristics. We focus on tasks including instruction-based molecule generation (Edwards et al. 2022) and molecule editing (Fang et al. 2023), striving to integrate two modalities of information in drug design: aligning text descriptions to ensure desired properties and generating graph structures that represent potential molecules.
Figure 1: A deep generative model that works like a drug design specialist, with instruction input and molecule output.
One widely adopted approach involves generating molecular strings from instructions, such as SMILES or SELFIES (Weininger 1988; Krenn et al. 2020), which serve as a stopgap for describing molecular graphs using language. Several recent studies have explored this method by employing auto-regressive language models  (Edwards et al. 2022; Liu et al. 2023d; Pei et al. 2023), deploying large language models  (Achiam et al. 2023; Ye et al. 2023; AI4Science and Quantum 2023), or modifying existing molecules based on specified instructions  (Liang et al. 2023).
However, despite these advancements, current methods face significant challenges. The primary limitation of molecular strings lies in their insufficient capability to accurately represent complex molecular structures, as it is widely recognized that they lack spatial structure information such as neighborhood (Jiang et al. 2022; Wu et al. 2023; Song et al. 2020), which could limit performance. Additionally, aligning the language space with the molecular space often results in deviations from natural molecular distributions, such as generating unstable or rare structures through simple string reassembly (Druchok et al. 2021). It is therefore essential to develop a multimodal model that possesses robust molecular structure representational capabilities while maintaining instruction-following abilities.
As a non-sequential data structure, graph generation capacity could be constrained by fixed orders imposed by auto-regressive models. This limitation motivates our usage of diffusion, a powerful non-sequential generative model. Therefore, we propose the Unified Text-Graph Diffusion Model (UTGDiff), a novel diffusion framework for text-guided graph generation. UTGDiff extends vanilla transformer as graph denoising networks through attention bias from graph edge categories, ensuring a strong capacity for learning graph structure representations and executing instruction texts without additional graph encoding modules. The equivariance of our model in graph structure ensures structural modeling abilities beyond asymmetric SMILES (O’Boyle and Dalke 2018; Arús-Pous et al. 2019).
We performed extensive experiments on the pre-trained UTGDiff model across diverse tasks. Our results show that UTGDiff achieves improved outcomes over language model baselines in drug design tasks (instruction-based molecule generation and molecule editing) despite its small parameter size (125M). Additionally, our use of a diffusion-generation pipeline and text-graph denoising network achieves a relatively high accuracy of over 85% in terms of chemical validity. In summary, our contributions are:
  • We present a unified text-graph diffusion framework for instruction-based graph generation, enabling the creation of molecular structures that are instructed by text and represent accurate chemical structures.
  • We propose UTGDiff, which advances the state of the art by employing this unified text-graph transformer that effectively captures the features of both graph and text data. The model distinguishes itself from prior work by enhancing molecular representation and generation.
  • Experimental results demonstrate that UTGDiff outperforms existing language model baselines in several critical areas of drug design, including instruction-based molecule generation and molecule editing, with an equivalent level of pretraining corpus.
Related Works
Molecule Generation
Early attempts at molecule generation introduced sequence-based models for SMILES strings (Dai et al. 2018; Gómez-Bombarelli et al. 2018). Due to the lack of structural information, recent studies have explored graph-based methods, including auto-regressive models (Grisoni et al. 2020; You et al. 2018), VAEs (Simonovsky and Komodakis 2018; Jin, Barzilay, and Jaakkola 2018; Liu et al. 2018), GANs (De Cao and Kipf 2018), and Normalizing Flows (Ma and Zhang 2021; Zang and Wang 2020). A relatively novel approach, offering superior performance and better scalability on larger molecules, is Graph Diffusion (Niu et al. 2020; Jo, Lee, and Hwang 2022; Vignac et al. 2022), which serves as our foundational framework.
Conditional Molecule Generation
Conditional molecule generation, an important topic in molecule generation, aims to generate a molecular graph that satisfies a specific property. Early attempts are based on a scalar property such as dipole moment $\mu$ and p-logP scores (the octanol-water partition coefficients) (Li, Zhang, and Liu 2018; Huang et al. 2023). More complex conditions involve textual instructions, with existing solutions including sequence-based generation (Edwards et al. 2022; Christofidellis et al. 2023; Liu et al. 2023d; Li et al. 2023; Pei et al. 2024) and multimodal generation employing a text encoder and graph generative network (Su et al. 2022; Liu et al. 2024).
Cross-Modal Molecule Models
In addition to the instruction-based cross-modal molecule generation models discussed above, some studies represent molecules and text in cross-modal models for molecule property prediction or text generation tasks (Liu et al. 2023c; Zhao et al. 2024; Liu et al. 2023a), employing contrastive alignment, cross-modal projector, or unified text-graph backbone.
Discrete Diffusion
Previous Gaussian noise in continuous diffusion models has been shown to be insufficient for discrete data structure. As a solution, researchers introduced the discrete diffusion method, initially discussed in text diffusion (Austin et al. 2021; He et al. 2022) and later adapted for graph data (Vignac et al. 2022; Kong et al. 2023).
Preliminaries
Figure 2: Overview of the UTGDiff framework. It incorporates attention bias into the vanilla transformer, forming a unified text-graph transformer that serves as a denoising network for discrete graph diffusion to generate molecular graphs from instructions. Noise decays some nodes and edges into [MASK] during training (forward process), with the reverse process aiming to recover original graphs as the training objective. Sampling starts with a masked graph and iterates $T$ times to reduce noise.
The objective of the instruction-based molecule generation task is to generate a molecular graph $G=(V,E)$ from a given textual description $S$ specifying molecular functions and properties. The graph consists of $m$ nodes $V\in\mathbb{R}^{m\times a}$, which represent node attributes, and edges $E\in\mathbb{R}^{m\times m\times b}$, which represent edge attributes. Each attribute $v_{i}\in\mathbb{R}^{a}$ and $e_{i,j}\in\mathbb{R}^{b}$ is a one-hot vector representing the categories of nodes and edges, respectively. The instruction text is defined as $S=[s_{1},\dots,s_{n}]$, where $s_{i}$ denotes the $i$-th token, and $n$ is the length of the instruction. The goal is to perform conditional graph generation $V,E\sim P(V,E\mid S)$.
Method
Overview of Diffusion Framework
We propose a text-conditioned diffusion framework for molecular graph generation, as depicted in Figure 2. As a generative model, the diffusion model iteratively refines random noise into structured data through a series of transformations. It involves the forward process, which gradually adds noise to data via a noise model, and the reverse process, which denoises the data to recover the original structure using a denoising network. At each reverse step, the model estimates and removes the noise component, progressively generating the final structure. The core component of our framework is the novel unified text-graph denoising network, described in the Reverse Process section.
Forward Process and Noise Model
The forward process, associated with the noise model $q$, progressively adds noise to corrupt the clean graph $G^{0}$, creating an increasingly noisy graph sequence $(G^{1},\dots,G^{T})$:
$q(G^{1:T}\mid G^{0})$ $=\prod^{T}_{t=1}q(G^{t}\mid G^{t-1})$ (1)
Since traditional Gaussian noise is not suitable for graphs due to its disruption of sparsity and graph theory properties such as connectivity, we follow the discrete diffusion setting in (Vignac et al. 2022). The noise model $q$ is represented by a sequence of transition matrices $(Q^{1},\dots,Q^{T})$. It acts on a one-hot encoding $x\in\mathbb{R}^{d}$ over $d$ categories, where $[Q^{t}]_{i,j}$ indicates the probability of transferring the category $i$ to $j$ at time step $t$. We independently diffuse each one-hot category-based attribution in nodes $V$ and edges $E$ using separate matrices $Q_{V}^{t}$ and $Q_{E}^{t}$. The forward process is:
$q(G^{t}$ $\mid G^{t-1})=(V^{t-1}Q_{V}^{t},E^{t-1}Q_{E}^{t})$ (2)
It can also be expressed as:
$q(G^{t}\mid G^{0})=(V\bar{Q}_{V}^{t},E\bar{Q}_{E}^{t})$ (3)
$\bar{Q}_{V}^{t}=Q_{V}^{1}\dots Q_{V}^{t},\quad\bar{Q}_{E}^{t}=Q_{E}^{1}\dots Q_{E}^{t}$
Specifically, we employ absorbing discrete diffusion (Austin et al. 2021), where each element $v_{i}^{t}$ (and $e_{i,j}^{t}$) independently decays into an absorbing state based on the noise schedule probabilities. The absorbing state is defined as a [MASK] token, shared by text and graph tokens. Following the noise schedule $\beta(t)=(T-t+1)^{-1}$, the corresponding transition matrix $Q^{t}$ is formally expressed as follows, applying to both nodes and edges simultaneously, where $z$ is the index for [MASK] token:
i,j =$\left\{\begin{array}[]{l l l}{{1}}&{{\mathrm{if}}}&{{i=j=z}}\\ {{1-\beta(t)}}&{{\mathrm{if}}}&{{i=j\not=z}}\\ {{\beta(t)}}&{{\mathrm{if}}}&{{j=z,i\not=z}}\end{array}\right.$ (4)
Reverse Process and Denoising Network
Overview
The reverse process relies on a denoising neural network $\phi_{\theta}$ parameterized by $\theta$. This network is trained to invert the forward process by predicting $G^{t-1}$ from $G^{t}$ using the text instruction $S$, formally represented as:
$p_{\theta}(G^{0:T-1}\mid G^{T},S)=\prod^{T}_{t=1}p_{\theta}(G^{t-1}\mid G^{t},S).$ (5)
As the formula indicates, the denoising network must simultaneously encode the textual instruction and graph to generate the probability distributions of nodes and edges at next time step. To ensure sufficient comprehension of instructions, we retain the transformer model and propose the Unified Text-Graph Transformer as our denoising network.
Tokenizer
To adapt the vanilla transformer for processing graph data, we first address the representation of graph nodes by treating them as new tokens. The initial embeddings are denoted as $h^{0}=\text{Emb}(S+V)=[h_{1},\dots,h_{n},h_{n+1},\dots,h_{n+m}]\in\mathbb{R}^{(n+m)\times d_{h}}$, corresponding to $m$ graph nodes and $n$ text tokens, where $h_{1},\dots,h_{n}$ are encodings derived from the token list $S=[s_{1},\dots,s_{n}]$ and $h_{n+1},\dots,h_{n+m}$ are encodings based on the categorical list of atoms in the molecule (e.g., [[S],[O],[O]] for sulfur oxide). Our model treats each graph category as a novel token to prevent interference from irrelevant tokens, thereby avoiding misrepresentations such as interpreting cobalt as a combination of carbon and oxygen.
Attention Bias
Next, we address the representation of edges. In the attention layer, which includes parameters for values, keys, queries, and outputs, i.e., $W_{V},W_{Q},W_{K}\in\mathbb{R}^{d_{h}\times d_{k}},W_{O}\in\mathbb{R}^{d_{k}\times d_{h}}$, we introduce an additional bias term to incorporate edge information. The attention score between the $i$-th and $j$-th graph tokens is now formalized as:
$\hat{A}_{i,j}^{l}$ $=\frac{1}{\sqrt{d_{k}}}\left(h_{i}^{l}W_{Q}\right)\left(h_{j}^{l}W_{K}\right)^{{}^{\prime}}+b_{i,j}^{l},$ (6)
$A^{l}$ $=\mathrm{softmax}(\hat{A}^{l}),\ \text{Attn}(H)=A^{l}H^{l}W_{V}W_{O}$
Here, $b_{i,j}$ represents the bias added to the graph tokens to incorporate structural information (Ying et al. 2021; Choromanski et al. 2022; Park et al. 2022; Zhao et al. 2023). The most fundamental requirement for encoding graph data is the ability to distinguish whether two nodes in the graph are neighbors, which corresponds to the expressive power of a single-layer GNN (Xu et al. 2018). In addition, edge attribution must maintain symmetry for structural interpretability. Representing the initial embedding from the edge category as $b_{i,j}^{0}$, we therefore define the attention bias as follows:
$b_{i,j}^{l}=\begin{cases}b_{i,j}^{0},&\text{if }l=0\\ A_{i,j}^{l-1},&\text{otherwise}\\ \end{cases}$ (7)
In addition to fulfilling the structural requirements outlined above and theoretically discussed in Appendix B, a vital advantage is that our model requires minimal modifications to existing language models, enabling them to process graph and text data simultaneously without additional graph neural network and benefiting from initial pre-trained weights. Moreover, our model enhances the integration of text and graph data, as the layer-level interactions between modalities facilitate a better alignment, allowing text information to be incorporated into graph data at multiple scales.
Output
After obtaining the final hidden states $h^{L}$ for tokens and $b^{L}$ for edges through stacked attention layers, we map these hidden states back to the vocabulary space to determine the probability of each category for diffusion. We employ two separate masked language model heads for nodes and edges at the end of the transformer to decode these hidden states to the distribution of a clean graph.
Ultimately, we obtain the nodes logits $\log p_{\theta}(v_{i}\mid G^{t},S)$ and edges logits $\log p_{\theta}(e_{i,j}\mid G^{t},S)$ for a clean graph as output. Assembling these logits into matrices as $\hat{p}(V)\in\mathbb{R}^{m\times a}$, $\hat{p}(E)\in\mathbb{R}^{m\times m\times b}$ , we summarize our unified text-graph denoising network as:
$\phi_{\theta}(G^{t},S)=\hat{p}(G)=[\hat{p}(V),\hat{p}(E)]$ (8)
Sampling
There remains a gap between the distribution of the clean graph and the required next-step distribution as formula (5). This gap is fixed through the $x_{0}$-parameterization, which adds noise back to the clean graph prediction, represented as $p_{\theta}(G^{t-1}\mid G^{t},S)=q(G^{t-1}\mid G^{t},\phi_{\theta}(G^{t},S))$, yielding a distribution from which $G^{t-1}$ is sampled:
$p_{\theta}(v^{t-1}_{i}\mid G^{t},S)=\sum_{v\in\mathcal{V}}q(v^{t-1}_{i}\mid v_{i}=v,v^{t}_{i})\;{\hat{p}}_{i}(v)$ (9)
$p_{\theta}(e^{t-1}_{i,j}\mid G^{t},S)=\sum_{e\in\mathcal{E}}q(e^{t-1}_{i,j}\mid e_{i,j}=e,e^{t}_{i,j})\;{\hat{p}}_{i,j}(e)$
Now sampling can be performed as Algorithm 2. It begins with graph $G^{T}$, where each node and edge is masked. At each step $t$, it calculates the distribution and samples the graph at step $t-1$, gradually reducing the noise and eventually resulting in a clean graph $G^{0}$. Our model can implicitly sample the molecular length through the empty token, which is crucial as target molecule may be unique for instructions.
Training Objective
Since the denoising model takes a noisy graph $G^{t}$ as input and aims to predict the clean graph $G^{0}$, we optimize the cross-entropy loss (represented as $\mathrm{CE}$) for each node and edge between the predicted probabilities $\hat{p}(G)$ and the clean graph $G$ during training. The diffusion loss is defined as:
loss $=l(\hat{p}(V),V)+l(\hat{p}(E),E)$ (10)
$=\sum_{1\leq i\leq m}\mathrm{CE}(v_{i},\hat{p}_{i}(V))+\sum_{1\leq i,j\leq m}\mathrm{CE}(e_{i,j},\hat{p}_{i,j}(E))$
In addition, as previously discussed, maintaining balanced training across different modalities is crucial to prevent the model from diminishing its text comprehension ability after extensive training, as the tokens of the graph and text are non-overlapping. To address this issue, we incorporate a textual loss component to preserve the model’s text comprehension capability, which is defined as the cross-entropy loss for each token between the predicted logits $\hat{p}(S)$ and the ground-truth token in the instruction. Through the denoising model, the logits can be directly obtained alongside $\hat{p}(V)$ following a masked language model head, expressed as $\phi_{\theta}(G^{t},S)=[\hat{p}(S),\hat{p}(V),\hat{p}(E)]$. The final loss is thus represented as:
$\text{loss}=l(\hat{p}(V),V)+l(\hat{p}(E),E)+l(\hat{p}(S),S)$ (11)
See Algorithm 1 for the complete training procedure.
Algorithm 1 Training Algorithm
0:  A graph $G=(V,E)$ with description $S$
1:  repeat
2:     Sample $t\sim U(1,\dots,T)$
3:     Sample $G^{t}\sim V\bar{Q}_{V}^{t}\times E\bar{Q}_{E}^{t}$
4:     $\hat{p}(S),\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
5:     loss = $l_{CE}(\hat{p}(E),E)+l_{CE}(\hat{p}(V),V)+l_{CE}(\hat{p}(S),S)$
6:     optimizer.step(loss)
7:  until converged
Algorithm 2 Sampling Algorithm
1:  $G^{T}$ with all masked nodes and edges
2:  for $t=T$ to $1$ do
3:     $\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
4:     $p_{\theta}(v^{t-1}_{i}\mid G^{t},S)=\sum_{v}q(v^{t-1}_{i}\mid v_{i}=v,v^{t}_{i})\;{\hat{p}}_{i}(v)$
5:     $p_{\theta}(e^{t-1}_{ij}\mid G^{t},S)=\sum_{e}q(e^{t-1}_{ij}\mid e_{ij}=e,e^{t}_{ij})\;{\hat{p}}_{ij}(e)$
6:     $G^{t-1}\sim\prod_{i}p_{\theta}(v^{t-1}_{i}\mid G^{t},S)\prod_{ij}p_{\theta}(e^{t-1}_{ij}\mid G^{t},S)$
7:  end for
8:  return  $G^{0}$
Type Model #params MACCS FTS $\uparrow$ RDK FTS $\uparrow$ Morgan FTS $\uparrow$ FCD $\downarrow$ Exact $\uparrow$ Valid $\uparrow$
Specialist Auto-regressive (w.o. pretrain) T5-small 77M 0.704 0.578 0.525 2.89 0.064 0.608
T5-base 248M 0.731 0.605 0.545 2.48 0.069 0.660
T5-large 783M 0.823 0.731 0.670 1.22 0.279 0.902
BioT5-base (reproduce) 252M 0.821 0.708 0.633 1.67 0.071 1.000
Specialist Auto-regressive (pretrain) MolT5-small 77M 0.703 0.568 0.517 2.49 0.079 0.721
MolT5-base 248M 0.721 0.588 0.529 2.18 0.081 0.772
MolT5-large 783M 0.834 0.746 0.684 1.20 0.311 0.905
MolXPT 350M 0.859 0.757 0.667 0.45 0.215 0.983
BioT5-base (reproduce) 252M 0.843 0.745 0.676 1.41 0.097 1.000
Text diffusion tgm-dlm 125M 0.854 0.739 0.688 0.77 0.242 0.871
tgm-dlm w/o corr 125M 0.874 0.771 0.722 0.89 0.242 0.789
LLM chatgpt3.5 (0-shot) 7B 0.703 0.568 0.517 2.49 0.079 0.721
Chatgpt3.5 (10-shot) 7B 0.847 0.708 0.624 0.57 0.139 0.887
Graph-based Momu-S 113M 0.244 0.103 0.047 22.21 0.000 1.000
UTGDiff (w.o. pretrain) 125M 0.867 0.763 0.695 0.92 0.227 0.856
UTGDiff (pretrain) 125M 0.885 0.795 0.724 0.86 0.374 0.893
Table 1: Results of the instruction-based molecule generation on ChEBI-20 dataset for both with/without pretraining setting.
Pretraining Method
Recent studies have demonstrated that BERT can be viewed as a one-step [MASK] absorbing diffusion model  (Devlin et al. 2019; Austin et al. 2021), since one-step noising replaces some tokens with [MASK] and denoising performs masked prediction. This shared training objective allows for the effective initialization from pre-trained language models (He et al. 2022). Therefore, to preserve language comprehension, we utilize the pre-trained RoBERTa model (Liu et al. 2019) for initialization.
However, the introduction of novel graph tokens during tokenization renders the direct loading of pre-trained BERT/RoBERTa parameters insufficient for mask prediction on graph tokens. Consequently, our pretraining methods are designed to equip the model with mask prediction capabilities for graph data while retaining its functionality for text data. We collect paired or single-modal data in both text and molecular graph modalities, mask all tokens and edge indices with a 15% probability, and pretrain the model using the masked language prediction method. The effectiveness of this pretraining approach for graph data has been validated by previous studies on graph pretraining (Hu et al. 2020; Hou et al. 2022).
Experiment
Dataset
Model names
molT5 C4, ZINC
BioT5 C4, Pubmed, ZINC (w./w.o. IUPAC),
pubchem324K, bioRxiv, NER Wrapped biotext
UTGDiff Pubmed, ZINC, pubchem324K
Table 2: Review of pretraining datasets. Some data listed in BioT5 is unavailable, and we reproduce it with our dataset.
Pretraining Dataset
We collect several open-source datasets for pretraining. For the molecule-text pair data, we select the PubChem-324K dataset (Liu et al. 2023c), which contains approximately 320,000 molecule-text pairs from PubChem. For the single-modal data, we utilize the molecular graph dataset ZINC20 (Irwin et al. 2020) and biomedical text data from PubMed abstracts (White 2020), totaling nearly 100 million entries. Due to computational budget constraints, we are unable to scale up the training size and batch size of the pretraining corpus to the extent achieved by BioT5 (Pei et al. 2023), which utilizes non-open-source pretraining data.
Generation Dataset
For molecule generation tasks, we utilize the ChEBI-20 dataset (Edwards, Zhai, and Ji 2021) for finetuning and evaluation, which consists of 33,010 entries, with 10% allocated for validation and 10% for testing. We adhere to the original data split settings.
Editing Dataset
For molecule editing, we select the retrosynthesis and forward reaction prediction tasks from Mol-Instructions (Fang et al. 2023). The retrosynthesis task begins with a target compound and identifies potential precursor molecules as outputs. The forward reaction prediction task involves forecasting likely products as output molecular graphs, given specific reactants and reagents. Each of these two datasets contains approximately 120,000 reactant-product pairs for training and 1,000 pairs for testing, with each entry accompanied by an instruction.
Figure 3: Some generation examples from our model and baseline models.
model 1-64 65-96 97-128 129-160 $>$ 160
BioT5 (reproduce) 0.810 0.866 0.859 0.817 0.788
UTGDiff 0.860 0.903 0.901 0.853 0.821
Table 3: Analysis for scalability on instruction length.
Baselines
Several prior studies have explored sequence-based models for evaluation. For the ChEBI-20 dataset, the baseline models include (1) specialist auto-regressive models such as T5 (Raffel et al. 2020), MolT5 (Edwards et al. 2022) and BioT5 (Pei et al. 2023); (2) LLMs such as ChatGPT-3.5 for zero-shot and 10-shot; and (3) text diffusion model tgm-dlm (Gong et al. 2024). For the editing datasets, the baseline models include (1) specialist auto-regressive models such as BioT5 (Pei et al. 2024) and TEXT+CHEM T5 (Christofidellis et al. 2023); (2) LLM such as InstructMol (Cao et al. 2023) and other models. Since BioT5’s pretraining data is not open-sourced, we reproduce their experiments training on our pretraining datasets using reported hyperparameters, rather than referencing their reported performance.
These baselines help us highlight the deficiencies of sequence-based models and clarify our improvements. Additionally, to assess the effectiveness of our diffusion method, we finetuned MoMu-S (Su et al. 2022) on the CHEBI-20 dataset for instruction-based generation. Due to the lack of capability to interpret language in the graph-only structure-aware network, there’re few baselines related to graph diffusion and it’s challenging to directly adapt them into these tasks. Therefore, we don’t consider them as baselines, but further discuss these methods in Ablation Study.
Type Model MACCS RDK Morgan Exact Valid
LLM LLAMA 0.029 0.018 0.017 0.000 0.010
GALACTICA 0.274 0.167 0.134 0.000 0.986
Mol-Instructions 0.487 0.283 0.230 0.009 1.000
Llama-7b* (LoRA) 0.294 0.136 0.106 0.000 1.000
InstructMol-G 0.523 0.422 0.285 0.114 1.000
InstructMol-GS 0.852 0.753 0.714 0.407 1.000
T5 TEXT+CHEM T5 0.765 0.685 0.585 0.141 0.698
BioT5+ (reproduce) 0.904 0.843 0.810 0.480 1.000
UTGDiff 0.904 0.847 0.817 0.541 0.945
Table 4: Mol-Instruction Retrosynthesis results.
Instruction-Based Molecule Generation
Results
Our results are shown in Table 1. The evaluation metrics are chosen based on previous studies, including a validity metric and five graph-level similarity metrics, with the exception of text-level similarity metrics such as BLEU. Here, we present the results of a single experiment. Detailed descriptions of these metrics, additional results and variance obtained from different seeds, and details for experiments such as hyperparams are available in the supplementary material C,D. Our experimental findings are as follows:
Our model outperforms all auto-regressive methods across almost all similarity metrics with the same level of pretraining corpus. UTGDiff exhibits substantial improvements in the Exact match score, with gains of at least 6% over the best auto-regressive baseline. Also, UTGDiff achieves the highest FTS score, surpassing baseline models by at least 2%. These results indicate the methodological advantages of diffusion paradigm, since it is better suited to capturing the complex structures inherent in molecules.
Compared to text diffusion models, UTGDiff also performs better in both similarity and validity metrics. Especially, UTGDiff exhibits a remarkable 15% improvement in the Exact match score. This indicates that the superior performance of our model is not solely attributable to the diffusion paradigm. Instead, by incorporating graph-based representations, our model leverages explicit structural information in graph data processing, which is not attainable with text diffusion methods.
Type Model MACCS RDK Morgan Exact Valid
LLM LLAMA 0.002 0.001 0.001 0.000 0.010
GALACTICA 0.127 0.036 0.051 0.000 0.995
Mol-Instructions 0.509 0.313 0.262 0.045 1.000
Llama-7b* (LoRA) 0.649 0.499 0.407 0.012 1.000
InstructMol-G 0.717 0.519 0.457 0.114 1.000
InstructMol-GS 0.878 0.776 0.741 0.407 1.000
T5 TEXT+CHEM T5 0.789 0.705 0.652 0.141 0.698
BioT5+ (reproduce) 0.954 0.907 0.890 0.684 1.000
UTGDiff 0.973 0.943 0.942 0.825 0.972
Table 5: Mol-Instruction forward reaction prediction results.
Our model is far better than the graph-based MoMu-S model. The low performance of MoMu-S may be attributed to the limitations of MoFlow, as it struggles to follow complex instructions and generate long molecules, showing the importance of unified diffusion.
Our model has fewer parameters than other models with comparable performance. Our model has only 125M parameter size, smaller than other T5 models and LLMs.
Our model performs marginally worse than the reported results of BioT5+ and the recently finetuned LLM (Li et al. 2024). This can be attributed to the larger pretraining corpora they used as discussed in the ablation study. Also, our model shows minor shortcomings in validity metrics due to inevitable valence-induced invalidity. The low value in FCD can be attributed to misleading influences from other generated molecules, as discussed in supplementary material A.
Case Study
We present several generated examples in Figure 3. The model successfully generates accurate results comparable to baseline models (Example 1). In some instances, it produces the correct solution when other models fail (Example 2). Also, it can effectively capture essential prompt words like ”tosylate” and generate the required sulfur atom, demonstrating its capability to produce contextually appropriate outputs (Example 3).
Scalability on Instruction
We also present an experiment assessing performance with varying complexities of instructions. We report the MACCS results for varying lengths of instructions and compare them with BioT5. The results indicate that our diffusion model achieves better outcomes across all intervals, demonstrating enhanced scalability when handling more complex instructions.
Instruction-Based Molecule Editing
We also conducted experiments on molecule editing tasks to demonstrate the model’s generalization capabilities across diverse tasks. Transferring from molecular generation to editing requires an additional source molecular graph as input, which is incorporated into the instruction input in graph format. The results are presented in Table 4 and Table 5.
[…]
UTGDiff 1 0.838 0.515 0.861
10 0.879 0.860 0.968
100 0.885 0.893 1.929
1000 0.881 0.901 11.388
BioT5 - 0.843 1.000 0.763
Table 6: Ablation study on diffusion framework. Steps = 1 is equal to without diffusion
[…]
w.o. unifying 0.583 28.47 0.753 -
3M-Diffusion - - - 0.871
UTGDiff 0.867 0.923 0.856 0.957
Table 7: Ablation study on unifying denoising network.
Our model outperforms baseline models in all fingerprint and exact match metrics, with gains of at least 2% in forward reaction prediction tasks. These results underscore the methodological advantages of our model over previous string-based auto-regressive models. When compared to InstructMol-G and InstructMol-GS, our model exhibits a remarkable doubling of the exact match score in forward reaction prediction, with at least a 5% improvement in all metrics across both tasks. These results highlight the significance of the diffusion generation method for these tasks, as InstructMol-G incorporates a 2D molecular graph encoder while retaining a sequence-based decoder approach.
Ablation Study
In our ablation study, we validate our model focusing on the impact of pretraining and model design. All ablation experiments were performed using the CHEBI-20 dataset.
Model Design
Our investigation of model design focuses on two key components to highlight the importance of novel modules: (a) Ablation of the diffusion module. Using the $x_{0}$-parameterization, we can perform inference with flexible step size during the reverse process, while the one-step logits represent a scenario without diffusion. The effectiveness of the diffusion module is demonstrated in Table 6. (b) Ablation of the unified denoising network. This experiment involves the independent use of the graph diffusion network and text encoder, guided by a similarity gradient (Liu et al. 2023b), without the unification of the network. Additionally, we draw from the recent 3M-Diffusion model (Zhu, Xiao, and Honavar 2024), a text-guided graph diffusion method that employs separate encoders, and evaluate our generation results using their Similarity metrics. The results, shown in Table 7, indicate the effectiveness of the unified network.
Model MACCS $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Train loss $\downarrow$ Entry
from scratch 0.867 0.923 0.856 1.58e-2 0
Pair-only 0.870 0.892 0.879 1.99e-3 300K
All 0.885 0.866 0.893 7.69e-4 100M
Table 8: Ablation study on pretrain data.
Pretraining
We conducted another ablation experiment to assess the impact of pretraining scale. The comparison, detailed in Table 8, highlights the critical roles of both single-modal and paired pretraining datasets and demonstrates the importance of scale, particularly concerning the validity.
Conclusions and Limitations
In this paper, we propose UTGDiff, an instruction-based framework that unifies graph and text data within a single transformer through attention bias, leveraging for the denoising generation of graph data. Our model is evaluated on tasks involving instruction-based molecule generation and editing, showing superior performance while requiring fewer parameters with the same level of pretraining corpus.
Several limitations constrain this research. A critical constraint is our inability to scale the pretraining corpus to a larger size, a factor that significantly benefits models like BioT5 and may have limited our model’s ability to generalize across different molecular structures. Additionally, further experiments on more advanced discrete diffusion frameworks should be explored.
Supplementary Material
A. Analysis of FCD
FCD is a metric originally used to measure the distance between two sets of distributions. Previous works employed this metric to compare the distance between the ground-truth test sets and generation sets in instruction-based molecule generation tasks, rather than comparing the similarity one by one as other metrics do.
Since this metric compares the distance between molecule distributions, it can be misled by molecules generated from other instructions. For example, FCD is equivariant under reordering, meaning it cannot distinguish errors where generated molecules and instructions are exactly cross-matched. For instance, the instruction $S_{1}$ match the molecule ‘CCO’, and instruction $S_{2}$ match the molecule ‘[He]’. If we generate ‘CC’ with instruction $S_{1}$ and ‘[He]’ with instruction $S_{2}$, there’s no doubt that FCD([‘CCO’, ‘[He]’], [‘CC’, ‘[He]’]) ¿ 0. However, If we generate ‘[He]’ with instruction $S_{1}$ and ‘CCO’ with instruction $S_{2}$, which is a totally irrelavant results, we’ll got an unexpected result FCD([‘CCO’, ‘[He]’], [‘[He]’, ‘CCO’]) = 0.
To demonstrate the effectiveness of our model in FCD, we divided the test set into 10 subsets and calculated the average and variance FCD across them. This method can partially reduce interference between different instructions. We include tgm-dlm for comparison, as it has a lower FCD, while MolXPT is not open-sourced.
Using this processing method, our model achieved better FCD results, as shown in Table 9. The difference in magnitude between the two settings is an inherent property of FCD, which is related to the number of molecule sets.
B. Proof of required property
The edge attribution must maintain symmetry. This is conducted since the start of embedding is symmetrical as they have same categories. And therefore the edge attribution maintains symmetry in each layer.
Another required property for graph generation model is to ensure that the node’s attribution in the graph is equivariant to reorderings, meaning that the matrices under reordering can represent the same graph. This property requires no positional embedding in the denoising network. However, we empirically found the importance of positional embedding for better performance.
We attribute this result to the rich position information in the text description, as shown in figure 4. So, the model will lose its perception of this essential information without position embedding. This issue is more significant in diffusion models since the located step must be later than the decoding of the atom.
Figure 4: An example of generation results w/w.o position embedding. Adding position embedding can locate the position of the iodo group.
Therefore, to ensure equivariance on the inference step while retaining the utilization of position embedding, We adopt the following strategy: We use position embedding as before during training and make a random permutation of the position index in sequence during inference. It can be proven that the model is still equivariant during generation, although it is trained to fit a specific order of decoding atoms.
The theorem to be proven in this section can be written as follows:
Theorem 1
(Equivariancy for Graph Generation) For any permutation $\pi$, the model will generate a graph with node feature $V$ and adjacency matrix $A$ with input position index $i$, satisfying $P_{i}(V,A)=P_{\pi^{T}i}(\pi^{T}V,\pi^{T}A\pi)$.
The proof aims to show that: If the model’s network architecture is equivariant with input mapping permutation and the training loss is permutation invariant, the model will generate a distribution reordering with position index.
Lemma 1
[…]
Proof 1
Since the defined loss function is the same for each node and edge, we have
$l(\pi.{\hat{G}},\pi.G)=\sum_{i}l_{X}(\pi.{\hat{X}}_{i},x_{\pi^{-1}(i)})$
$+\sum_{i,j}l_{E}(\pi.{\hat{\mathbb{E}}}_{i,j},e_{\pi^{-1}(i),\pi^{-1}(j)})$
$=\sum_{i}l_{X}({\hat{X}}_{i},x_{i})+\sum_{i,j}l_{E}({\hat{\mathbb{E}}}_{i,j},e_{i,j})$
$=l(\hat{G},G)$
Lemma 2
[…]
Proof 2
Define $G_{t}=(X_{t},E_{t})$ as the noisy graph, $(\pi.X_{t},\pi.E_{t})$ as the permutation. Since the input is permutation equivariant, and we have:
  • The self attention architecture is permutation invariant
  • The linear layers are permutation invariant.
  • The Layer normalization is permutation equivariant.
Therefore, the model is a combination of permutation equivariant blocks.
Proof 3
proof of theorem 1
For the input embedding after model permutation, we have $\pi^{T}X+\pi^{T}emb(i)=\pi^{T}(X+emb(i))$, Therefore, the assumption of input mapping permutation equivariant holds.
Following the proof in DiGress, with the lemma(Xu et al. 2022): if a distribution $p(G_{T})$ is invariant to the action of a group G and the transition probabilities $p(G_{t-1}|G_{t})$ are equivariant, them $p(G_{0})$is invariant to the action of G. We apply this result to the special case of permutations:
  • The initial noise distribution is the mask distribution on each node and edge. It is therefore permutation invariant.
  • The denoising neural networks is permutation equivariant.
  • The transition probabilities function $p_{\theta}(G_{t-1}|G_{t})=\sum_{G}q(G_{t-1},G|G_{t})\hat{p}_{\theta}(G)$ is equivariant to $\bar{p}_{\theta}(G)$ and $G_{T}$.
The conditions are therefore satisfied, and the model satisfies $P_{i}(V,A)=P_{\pi^{T}i}(\pi^{T}V,\pi^{T}A\pi)$.
Model 10 subset (new) 1 set (origin)
Tgm-dlm $3.38\pm 0.078$ 0.77
UTGDiff $3.25\pm 0.055$ 0.866
Table 9: FCD comparison
C. Details for evaluation metrics
Since the model generates a matrix of nodes and edges, although ultimately a SMILES will be parsed based on the matrix, it is not directly generated, which can easily cause the problem of different strings for same graph. Therefore, although molecules can be represented by biological sequence structures, and previous models have also established evaluation indicators from the level of string similarity, this article does not involve these indicators in comparison, including NLP indicators such as BLEU(Papineni et al. 2002), Levenshtein(Miller, Vandome, and McBrewster 2009).
Therefore, we use a series of metrics related to the similarity of molecular graphs:
  • Exact: Whether the two molecules are same.
  • Valid: Whether the generated molecule satisfied the constraint for molecule, such as the valence rule.
  • FTS: We employ three fingerprint metrics: MACCS FTS, RDK FTS, and Morgan FTS, where FTS stands for fingerprint Tanimoto similarity. MACCS (Durant et al. 2002), RDK(Schneider, Sayle, and Landrum 2015) and Morgan(Rogers and Hahn 2010). The fingerprints of two molecules are compared using Tanimoto similarity (also known as Jaccard index), and the average similarity over the evaluation dataset is reported. We use RDKIT toolkit.
  • FCD score (Fréchet chemnet distance)(Preuer et al. 2018): Measure molecular similarity based on a pre-trained ”ChemNet” bioinformatics network. We use fcd 1.1 in python.
D. Details for experiments
Here’re the hyperparam for three different tasks describe in article, and here’s some general information on hyperparam:
(1) The pretraining process spans nearly 300K steps and is executed on four NVIDIA 24GB GeForce RTX 3090 GPUs with batch size 32 per GPU, totally trained for near 1 week. The finetuning process spans nearly 800K steps and is executed on two NVIDIA 24GB GeForce RTX 3090 GPUs with batch size 16 per GPU, totally trained for near 3 days.
(2) During finetuning, we gradually increase the accumulation step for the tradeoff of training efficiency and convergence level. The effective of this method is shown in the ablation of accumulation step shown in Table 10. So, the initial accumulation step is 1, and it will be increase to 64 finally.
(3) When the model is used for generation, we additionally introduces top-k sampling to help improve the quality of generation: when a class is taken from the calculated probability for discrete generation, only the larger node class is taken.
(4) All the training and generation program is running under specific seed. We only generate onces, but experiments shows there’s no significant variance during sampling for different seeds. We show the 3 seed example in forward reaction prediction in Table 14
The specific data is listed in Table 11, 12, 13
Prompt
Here we list the prompt for the three datasets: For CHEBI-20 dataset, We give its instruction form as : ”’[molecule description]’ is the description of molecule:” For the two editing dataset, since there’s task instruction in datasets, we don’t use any additional instructions. Here’s an example in Retrosynthesis dataset: ”Please suggest potential reactants for the given product.”; Here’s an example in forward reaction prediction dataset: ”With the provided reactants and reagents, propose a potential product.”
Model MACCS $\uparrow$ RDK $\uparrow$ Morgan $\uparrow$ FCD $\downarrow$ Exact $\uparrow$ Valid $\uparrow$
baseline 0.867 0.763 0.695 0.923 0.227 0.856
acc_step = 4 0.838 0.719 0.629 1.651 0.149 0.792
Table 10: ablation of accumulation step (from-scratch)
param name value
learning rate 5e-5
batch size 16
accumulation step (1,4,16,64)
accumulation update epoch (1,4,16,64)
top k 15
predict molecule length 128
seed 42
step size 20
Table 11: Hyperparam for CHEBI-20 datasets
param name value
learning rate 5e-5
batch size 16
accumulation step (1,4,16,64)
epoch 1000
accumulation update epoch (90,150,180)
top k 15
predict molecule length 108
seed 42
step size 10
Table 12: Hyperparam for Retrosynthesis datasets
param name value
learning rate 5e-5
batch size 16
accumulation step (1,4,16,64)
epoch 210
accumulation update epoch (90,150,180)
top k 15
predict molecule length 96
seed 42
step size 10
Table 13: Hyperparam for forward reaction prediction datasets
Type Model MACCS RDK Morgan Exact Valid
seed 0 0.971 0.938 0.938 0.821 0.973
seed 1 0.974 0.940 0.938 0.825 0.968
seed 42 0.973 0.943 0.942 0.825 0.972
Table 14: 3 seeds for forward reaction prediction results
Figure 5: The training curve in first 300K steps. Pretraining also demonstrates lower initial loss and faster convergence compared to training models from scratch.
E. Source of datasets
All the datasets we used are open-sorcued, can be founded in github or huggingface:
huggingface: ZINC(zpn/zinc20); Pubmed; Pubchem
github: Mol-instruction, CHEBI-20

### v2

@@
 Abstract
-Recent advancements in computational chemistry have increasingly focused on synthesizing molecules based on textual instructions.
-Integrating graph generation with these instructions is complex, leading most current methods to use molecular sequences with pre-trained large language models.
-In response to this challenge, we propose a novel framework, named UTGDiff (Unified Text-Graph Diffusion Model), which utilizes language models for discrete graph diffusion to generate molecular graphs from instructions.
-UTGDiff features a unified text-graph transformer as the denoising network, derived from pre-trained language models and minimally modified to process graph data through attention bias.
-Our experimental results demonstrate that UTGDiff consistently outperforms sequence-based baselines in tasks involving instruction-based molecule generation and editing, achieving superior performance with fewer parameters given an equivalent level of pretraining corpus.
-Our code is availble at https://github.com/ran1812/UTGDiff.
-Introduction
-Molecules, as fundamental units of matter, possess structures and properties that determine the characteristics of substances.
-Research into novel molecules is pivotal for chemistry, biology, and drug discovery (Schneider and Fechner 2005; Hughes et al.
-2011), as these fields depend on the discovery of molecules with specific properties.
-A significant challenge in drug discovery is designing molecules that exhibit desired features, maintain structural stability, and interact with specific targets (Hartenfeller and Schneider 2011).
-Traditionally, this process has been resource-intensive and time-consuming (Dickson and Gagnon 2009; Khanna 2012).
-However, recent advancements in deep learning have revolutionized drug design by offering cost-effective methods (Nag et al.
-2022; Askr et al.
-2023), drawing significant attention to the research of generating drug-like molecules (Bagal et al.
-2021; Schneuing et al.
-2022).
-In this work, we aim to bridge the gap between molecular structures and natural language, designing a generative model that functions akin to a drug design expert (illustrated in Figure 1).
-By receiving specific instructions, this model can precisely generate candidate molecules with the indicated characteristics.
-We focus on tasks including instruction-based molecule generation (Edwards et al.
-2022) and molecule editing (Fang et al.
-2023), striving to integrate two modalities of information in drug design: aligning text descriptions to ensure desired properties and generating graph structures that represent potential molecules.
-Figure 1: A deep generative model that works like a drug design specialist, with instruction input and molecule output.
-One widely adopted approach involves generating molecular strings from instructions, such as SMILES or SELFIES (Weininger 1988; Krenn et al.
-2020), which serve as a stopgap for describing molecular graphs using language.
-Several recent studies have explored this method by employing auto-regressive language models  (Edwards et al.
-2022; Liu et al.
-2023d; Pei et al.
-2023), deploying large language models  (Achiam et al.
-2023; Ye et al.
-2023; AI4Science and Quantum 2023), or modifying existing molecules based on specified instructions  (Liang et al.
-2023).
-However, despite these advancements, current methods face significant challenges.
-The primary limitation of molecular strings lies in their insufficient capability to accurately represent complex molecular structures, as it is widely recognized that they lack spatial structure information such as neighborhood (Jiang et al.
-2022; Wu et al.
-2023; Song et al.
-2020), which could limit performance.
-Additionally, aligning the language space with the molecular space often results in deviations from natural molecular distributions, such as generating unstable or rare structures through simple string reassembly (Druchok et al.
-2021).
-It is therefore essential to develop a multimodal model that possesses robust molecular structure representational capabilities while maintaining instruction-following abilities.
-As a non-sequential data structure, graph generation capacity could be constrained by fixed orders imposed by auto-regressive models.
-This limitation motivates our usage of diffusion, a powerful non-sequential generative model.
-Therefore, we propose the Unified Text-Graph Diffusion Model (UTGDiff), a novel diffusion framework for text-guided graph generation.
-UTGDiff extends vanilla transformer as graph denoising networks through attention bias from graph edge categories, ensuring a strong capacity for learning graph structure representations and executing instruction texts without additional graph encoding modules.
-The equivariance of our model in graph structure ensures structural modeling abilities beyond asymmetric SMILES (O’Boyle and Dalke 2018; Arús-Pous et al.
-2019).
-We performed extensive experiments on the pre-trained UTGDiff model across diverse tasks.
-Our results show that UTGDiff achieves improved outcomes over language model baselines in drug design tasks (instruction-based molecule generation and molecule editing) despite its small parameter size (125M).
-Additionally, our use of a diffusion-generation pipeline and text-graph denoising network achieves a relatively high accuracy of over 85% in terms of chemical validity.
-In summary, our contributions are:
-  • We present a unified text-graph diffusion framework for instruction-based graph generation, enabling the creation of molecular structures that are instructed by text and represent accurate chemical structures.
-  • We propose UTGDiff, which advances the state of the art by employing this unified text-graph transformer that effectively captures the features of both graph and text data.
-The model distinguishes itself from prior work by enhancing molecular representation and generation.
-  • Experimental results demonstrate that UTGDiff outperforms existing language model baselines in several critical areas of drug design, including instruction-based molecule generation and molecule editing, with an equivalent level of pretraining corpus.
-Related Works
-Molecule Generation
-Early attempts at molecule generation introduced sequence-based models for SMILES strings (Dai et al.
-2018; Gómez-Bombarelli et al.
-2018).
-Due to the lack of structural information, recent studies have explored graph-based methods, including auto-regressive models (Grisoni et al.
-2020; You et al.
-2018), VAEs (Simonovsky and Komodakis 2018; Jin, Barzilay, and Jaakkola 2018; Liu et al.
-2018), GANs (De Cao and Kipf 2018), and Normalizing Flows (Ma and Zhang 2021; Zang and Wang 2020).
-A relatively novel approach, offering superior performance and better scalability on larger molecules, is Graph Diffusion (Niu et al.
-2020; Jo, Lee, and Hwang 2022; Vignac et al.
-2022), which serves as our foundational framework.
-Conditional Molecule Generation
-Conditional molecule generation, an important topic in molecule generation, aims to generate a molecular graph that satisfies a specific property.
-Early attempts are based on a scalar property such as dipole moment $\mu$ and p-logP scores (the octanol-water partition coefficients) (Li, Zhang, and Liu 2018; Huang et al.
-2023).
-More complex conditions involve textual instructions, with existing solutions including sequence-based generation (Edwards et al.
-2022; Christofidellis et al.
-2023; Liu et al.
-2023d; Li et al.
-2023; Pei et al.
-2024) and multimodal generation employing a text encoder and graph generative network (Su et al.
-2022; Liu et al.
-2024).
-Cross-Modal Molecule Models
-In addition to the instruction-based cross-modal molecule generation models discussed above, some studies represent molecules and text in cross-modal models for molecule property prediction or text generation tasks (Liu et al.
-2023c; Zhao et al.
-2024; Liu et al.
-2023a), employing contrastive alignment, cross-modal projector, or unified text-graph backbone.
-Discrete Diffusion
-Previous Gaussian noise in continuous diffusion models has been shown to be insufficient for discrete data structure.
-As a solution, researchers introduced the discrete diffusion method, initially discussed in text diffusion (Austin et al.
-2021; He et al.
-2022) and later adapted for graph data (Vignac et al.
-2022; Kong et al.
-2023).
-Preliminaries
+Recent advancements in computational chemistry have increasingly emphasized generating and editing molecules from textual instructions.
+However, integrating graph generation with instruction understanding remains challenging, as most existing approaches either rely on molecular sequences in text modality with limited structural information, or struggle with multimodal alignment in graph diffusion methods.
+To address these limitations, we propose UTGDiff (Unified Text-Graph Diffusion Model), a novel framework that utilizes pre-trained language models for discrete graph diffusion, enabling the generation of molecular graphs from instructions.
+UTGDiff introduces a unified text-graph transformer as a denoising network, adapted with minimal modifications from language models to process graph data via attention bias.
+Experimental results show that UTGDiff consistently outperforms both sequence-based and conditional graph-diffusion baselines on instruction-based molecule generation and editing tasks with fewer parameters, covering instructions specifying molecular structures or properties.
+[…]
+4522
+1 Introduction
+Molecules possess structures and properties that determine the characteristics of substances.
+Research into novel molecules is crucial for fields such as chemistry, biology, and drug discovery [43].
+A central challenge in drug discovery is designing molecules that are stable, exhibit desired functionalities, and interact with specific biological targets [17].
+Traditionally, this process has been resource-intensive and time-consuming [8].
+However, recent advances in deep learning have enabled cost-effective methods, revolutionizing drug design and accelerating the generation of drug-like molecules [38, 1].
+Modern drug discovery requires exploring vast chemical spaces under specific constraints to identify molecules that fulfill complex therapeutic objectives.
+Traditional approaches based on a few scalar properties (e.g., solubility) often fall short in capturing the diversity and complexity of real-world design targets, such as specific structural motifs or functional groups.
+In contrast, natural language offers a flexible and expressive medium for conveying high-level design intents that encompass functional and structural goals.
+Consequently, there is increasing interest in the tasks of instruction-based molecule generation and editing [12, 13, 31], which aims to generate novel candidates from natural language instructions or edit existing ones to meet desired instructions, as illustrated in Figure 1.
+Figure 1: A generative model works like a drug design specialist, with instruction input and molecule output.
+To generate molecules from natural language instructions, prior studies directly use language models on string-based representations like SMILES or SELFIES [48, 27].
+This method has been explored by fine-tuning small-scale language models like T5 [12] or adapting large language models [53].
+However, such single-modal methods face inherent limitations.
+Notably, molecular strings lack explicit structural information, such as neighborhood relationships based on edges between atoms [23, 51].
+Moreover, representing molecules purely in language can lead to deviation from realistic molecular distributions, as minor string reassembly may produce chemically implausible structures [9].
+These issues highlight the need for a multimodal framework for language and graph models, combining robust structural representation with strong instruction-following abilities.
+To better represent graph structures in multimodal frameworks, several recent studies have explored powerful graph generation techniques.
+Among them, graph diffusion methods have shown strong performance and scalability on larger molecules [40, 25].
+Recent efforts have introduced guiding the diffusion process using several scalar conditions via predictor-guided [47, 49] or predictor-free approaches [31, 59].
+However, when conditioning on natural language, existing methods face notable limitations.
+While sentences embeddings can replace scalar conditions, they are typically encoded as fixed vectors and injected into node and edge embeddings through simple operations like addition or linear projection.
+This prevents graph elements from attending to relevant textual tokens, limiting semantic selectivity and hindering fine-grained injection from text, as nodes and edges cannot determine which instruction components are relevant to themselves.
+Moreover, language representations remain static throughout denoising, preventing interaction and evolving with graph, thereby limiting the model’s ability to refine textual representations and provide guidance tailored to incomplete structures in denoising.
+These limitations highlight the need for finer-grained and bidirectional integration between language and graph.
 Figure 2: Overview of the UTGDiff framework.
@@
 Noise decays some nodes and edges into [MASK] during training (forward process), with the reverse process aiming to recover original graphs as the training objective.
-Sampling starts with a masked graph and iterates $T$ times to reduce noise.
-The objective of the instruction-based molecule generation task is to generate a molecular graph $G=(V,E)$ from a given textual description $S$ specifying molecular functions and properties.
-The graph consists of $m$ nodes $V\in\mathbb{R}^{m\times a}$, which represent node attributes, and edges $E\in\mathbb{R}^{m\times m\times b}$, which represent edge attributes.
-Each attribute $v_{i}\in\mathbb{R}^{a}$ and $e_{i,j}\in\mathbb{R}^{b}$ is a one-hot vector representing the categories of nodes and edges, respectively.
-The instruction text is defined as $S=[s_{1},\dots,s_{n}]$, where $s_{i}$ denotes the $i$-th token, and $n$ is the length of the instruction.
-The goal is to perform conditional graph generation $V,E\sim P(V,E\mid S)$.
-Method
-Overview of Diffusion Framework
-We propose a text-conditioned diffusion framework for molecular graph generation, as depicted in Figure 2.
-As a generative model, the diffusion model iteratively refines random noise into structured data through a series of transformations.
-It involves the forward process, which gradually adds noise to data via a noise model, and the reverse process, which denoises the data to recover the original structure using a denoising network.
-At each reverse step, the model estimates and removes the noise component, progressively generating the final structure.
-The core component of our framework is the novel unified text-graph denoising network, described in the Reverse Process section.
-Forward Process and Noise Model
-The forward process, associated with the noise model $q$, progressively adds noise to corrupt the clean graph $G^{0}$, creating an increasingly noisy graph sequence $(G^{1},\dots,G^{T})$:
-$q(G^{1:T}\mid G^{0})$ $=\prod^{T}_{t=1}q(G^{t}\mid G^{t-1})$ (1)
-Since traditional Gaussian noise is not suitable for graphs due to its disruption of sparsity and graph theory properties such as connectivity, we follow the discrete diffusion setting in (Vignac et al.
-2022).
-The noise model $q$ is represented by a sequence of transition matrices $(Q^{1},\dots,Q^{T})$.
-It acts on a one-hot encoding $x\in\mathbb{R}^{d}$ over $d$ categories, where $[Q^{t}]_{i,j}$ indicates the probability of transferring the category $i$ to $j$ at time step $t$.
-We independently diffuse each one-hot category-based attribution in nodes $V$ and edges $E$ using separate matrices $Q_{V}^{t}$ and $Q_{E}^{t}$.
-The forward process is:
+Sampling starts with a masked graph and iterates $T$ times to reduce noise, where each time it process together with fixed given instruction.
+To this end, we propose the Unified Text-Graph Diffusion Model (UTGDiff), a novel conditional graph diffusion framework that unifies language and graph modalities within a single language transformer.
+By embedding both modalities into a unified transformer, UTGDiff enables more effective token-level text injection, as graph tokens can selectively attend to text tokens via attention.
+In addition, the instruction representation is continuously updated through attention with graph tokens, allowing it to evolve with molecular structures and contribute back to graph generation.
+UTGDiff extends masked language models to simulate graph diffusion through three key novel modifications, treating [MASK] tokens as noise in graph, extending the vocabulary with atom-specific tokens, and introducing a loss function to preserve text comprehension for alignment.
+We further incorporate attention bias into the transformer to model edges.
+We conduct extensive experiments on the pre-trained UTGDiff model across instruction-based molecule generation and editing tasks, covering conditions on both molecular structures and properties.
+Our results demonstrate that, despite its small parameter size, UTGDiff consistently outperforms both language model and graph diffusion baselines under comparable pretraining resources available to us, achieving higher similarity while maintaining a relatively high chemical validity.
+In summary, our main contributions are:
+  • We propose UTGDiff, a novel framework that unifies language and graph modalities within a single language transformer, with graph selectively attending to text and text being refined by graph.
+  • We extend language models to simulate graph diffusion by treating [MASK] as graph noise, extending atom-specific tokens, modifying the loss function, and incorporating attention bias for edges.
+  • We conduct extensive experiments showing that UTGDiff outperforms language model and graph diffusion baselines in key areas of drug design, achieving higher similarity and relatively high validity under comparable pretraining resources available to us.
+2 Related Works
+Molecule Generation
+Early works on molecule generation introduced sequence-based models for SMILES [6, 14].
+Due to the lack of structural information, recent studies explored graph-based methods, including auto-regressive [16, 55], VAEs [45, 24], GANs [7], and Normalizing Flows [36, 56].
+Graph Diffusion is a relatively new approach with superior performance [25, 47], which serves as our foundational framework.
+Conditional Molecule Generation
+Conditional molecule generation aims to generate a molecule that satisfies a specific property.
+Early attempts are based on single scalar property such as dipole moment $\mu$ and p-logP scores [30, 21].
+More complex conditions involve multi-condition [3, 31] and textual instructions.
+The latter are addressed by sequence-based generation [12, 5, 28, 41] and multimodal generation employing a graph decoder [46, 32].
+Cross-Modal Molecule Models
+In addition to the above instruction-based cross-modal graph generation models, some studies represent molecules and text in cross-modal models for property prediction or text generation [35, 58, 33], employing contrastive alignment, cross-modal projector, or unified text-graph backbone.
+Discrete Diffusion
+Previous Gaussian noise in continuous diffusion models has been shown to be insufficient for discrete data structure.
+As a solution, researchers introduced the discrete diffusion method, initially discussed in text diffusion [2, 18] and later adapted for graph data [47, 26].
+3 Method
+3.1 Problem Formulation
+The objective of our task is to generate a molecular graph $G=(V,E)$ from a given textual description $S$ specifying molecular functions and properties.
+The graph consists of $m$ nodes $V\in\mathbb{R}^{m\times a}$ and edges $E\in\mathbb{R}^{m\times m\times b}$ representing node and edge attributes, respectively.
+Each attribute $v_{i}\in\mathbb{R}^{a}$ and $e_{i,j}\in\mathbb{R}^{b}$ is a one-hot vector corresponding to the categories of the respective nodes and edges.
+The instruction text is defined as $S=[s_{1},\dots,s_{n}]$, where $s_{i}$ denotes the $i$-th token and $n$ is its length.
+The goal is to perform conditional graph generation, i.e., sampling $V,E\sim P(V,E\mid S)$.
+3.2 Overview of Diffusion Framework
+Building upon the above problem setting, We propose a text-conditioned diffusion framework for molecular graph generation, illustrated in Figure 2.
+As a generative process, the diffusion model iteratively refines random noise into structured data through a sequence of transformations.
+It involves a forward process which progressively corrupts the data by adding noise according to a noise model, and a reverse process which progressively denoises the data to reconstruct the original molecular structure using a denoising network.
+The core component of our framework is the unified text-graph denoising transformer, started from a language-pretrained transformer for text comprehension, and enhanced with a [MASK] noise module and graph-specific modeling techniques.
+3.3 Forward Process and Noise Model
+We adopt the forward diffusion process following Ho et al.
+[19], where the clean graph $G^{0}$ is progressively corrupted by a noise model $q$, producing a sequence of increasingly noisy graphs $(G^{1},\dots,G^{T})$.
+It is defined as $q(G^{1:T}\mid G^{0})=\prod^{T}_{t=1}q(G^{t}\mid G^{t-1})$.
+Following the discrete diffusion setting in [47], a sequence of transition matrices $(Q^{1},\dots,Q^{T})$ is defined independently for nodes $V$ and edges $E$ as noise model, denoted as $Q_{V}^{t}$ and $Q_{E}^{t}$.
+Each step in forward process can be expressed as:
 $q(G^{t}\mid G^{t-1})=(V^{t-1}Q_{V}^{t},E^{t-1}Q_{E}^{t})$ (1)
-It can also be expressed as:
-$q(G^{t}\mid G^{0})=(V\bar{Q}_{V}^{t},E\bar{Q}_{E}^{t})$ (3)
-$\bar{Q}_{V}^{t}=Q_{V}^{1}\dots Q_{V}^{t},\quad\bar{Q}_{E}^{t}=Q_{E}^{1}\dots Q_{E}^{t}$
-Specifically, we employ absorbing discrete diffusion (Austin et al.
-2021), where each element $v_{i}^{t}$ (and $e_{i,j}^{t}$) independently decays into an absorbing state based on the noise schedule probabilities.
-The absorbing state is defined as a [MASK] token, shared by text and graph tokens.
-Following the noise schedule $\beta(t)=(T-t+1)^{-1}$, the corresponding transition matrix $Q^{t}$ is formally expressed as follows, applying to both nodes and edges simultaneously, where $z$ is the index for [MASK] token:
-i,j =$\left\{\begin{array}[]{l l l}{{1}}&{{\mathrm{if}}}&{{i=j=z}}\\ {{1-\beta(t)}}&{{\mathrm{if}}}&{{i=j\not=z}}\\ {{\beta(t)}}&{{\mathrm{if}}}&{{j=z,i\not=z}}\end{array}\right.$ (4)
-Reverse Process and Denoising Network
-Overview
-The reverse process relies on a denoising neural network $\phi_{\theta}$ parameterized by $\theta$.
-This network is trained to invert the forward process by predicting $G^{t-1}$ from $G^{t}$ using the text instruction $S$, formally represented as:
-$p_{\theta}(G^{0:T-1}\mid G^{T},S)=\prod^{T}_{t=1}p_{\theta}(G^{t-1}\mid G^{t},S).$ (5)
-As the formula indicates, the denoising network must simultaneously encode the textual instruction and graph to generate the probability distributions of nodes and edges at next time step.
-To ensure sufficient comprehension of instructions, we retain the transformer model and propose the Unified Text-Graph Transformer as our denoising network.
-Tokenizer
-To adapt the vanilla transformer for processing graph data, we first address the representation of graph nodes by treating them as new tokens.
-The initial embeddings are denoted as $h^{0}=\text{Emb}(S+V)=[h_{1},\dots,h_{n},h_{n+1},\dots,h_{n+m}]\in\mathbb{R}^{(n+m)\times d_{h}}$, corresponding to $m$ graph nodes and $n$ text tokens, where $h_{1},\dots,h_{n}$ are encodings derived from the token list $S=[s_{1},\dots,s_{n}]$ and $h_{n+1},\dots,h_{n+m}$ are encodings based on the categorical list of atoms in the molecule (e.g., [[S],[O],[O]] for sulfur oxide).
-Our model treats each graph category as a novel token to prevent interference from irrelevant tokens, thereby avoiding misrepresentations such as interpreting cobalt as a combination of carbon and oxygen.
-Attention Bias
-Next, we address the representation of edges.
-In the attention layer, which includes parameters for values, keys, queries, and outputs, i.e., $W_{V},W_{Q},W_{K}\in\mathbb{R}^{d_{h}\times d_{k}},W_{O}\in\mathbb{R}^{d_{k}\times d_{h}}$, we introduce an additional bias term to incorporate edge information.
-The attention score between the $i$-th and $j$-th graph tokens is now formalized as:
-$\hat{A}_{i,j}^{l}$ $=\frac{1}{\sqrt{d_{k}}}\left(h_{i}^{l}W_{Q}\right)\left(h_{j}^{l}W_{K}\right)^{{}^{\prime}}+b_{i,j}^{l},$ (6)
-$A^{l}$ $=\mathrm{softmax}(\hat{A}^{l}),\ \text{Attn}(H)=A^{l}H^{l}W_{V}W_{O}$
-Here, $b_{i,j}$ represents the bias added to the graph tokens to incorporate structural information (Ying et al.
-2021; Choromanski et al.
-2022; Park et al.
-2022; Zhao et al.
-2023).
-The most fundamental requirement for encoding graph data is the ability to distinguish whether two nodes in the graph are neighbors, which corresponds to the expressive power of a single-layer GNN (Xu et al.
-2018).
-In addition, edge attribution must maintain symmetry for structural interpretability.
-Representing the initial embedding from the edge category as $b_{i,j}^{0}$, we therefore define the attention bias as follows:
-$b_{i,j}^{l}=\begin{cases}b_{i,j}^{0},&\text{if }l=0\\ A_{i,j}^{l-1},&\text{otherwise}\\ \end{cases}$ (7)
-In addition to fulfilling the structural requirements outlined above and theoretically discussed in Appendix B, a vital advantage is that our model requires minimal modifications to existing language models, enabling them to process graph and text data simultaneously without additional graph neural network and benefiting from initial pre-trained weights.
-Moreover, our model enhances the integration of text and graph data, as the layer-level interactions between modalities facilitate a better alignment, allowing text information to be incorporated into graph data at multiple scales.
-Output
-After obtaining the final hidden states $h^{L}$ for tokens and $b^{L}$ for edges through stacked attention layers, we map these hidden states back to the vocabulary space to determine the probability of each category for diffusion.
-We employ two separate masked language model heads for nodes and edges at the end of the transformer to decode these hidden states to the distribution of a clean graph.
-Ultimately, we obtain the nodes logits $\log p_{\theta}(v_{i}\mid G^{t},S)$ and edges logits $\log p_{\theta}(e_{i,j}\mid G^{t},S)$ for a clean graph as output.
-Assembling these logits into matrices as $\hat{p}(V)\in\mathbb{R}^{m\times a}$, $\hat{p}(E)\in\mathbb{R}^{m\times m\times b}$ , we summarize our unified text-graph denoising network as:
-$\phi_{\theta}(G^{t},S)=\hat{p}(G)=[\hat{p}(V),\hat{p}(E)]$ (8)
-Sampling
-There remains a gap between the distribution of the clean graph and the required next-step distribution as formula (5).
-This gap is fixed through the $x_{0}$-parameterization, which adds noise back to the clean graph prediction, represented as $p_{\theta}(G^{t-1}\mid G^{t},S)=q(G^{t-1}\mid G^{t},\phi_{\theta}(G^{t},S))$, yielding a distribution from which $G^{t-1}$ is sampled:
-$p_{\theta}(v^{t-1}_{i}\mid G^{t},S)=\sum_{v\in\mathcal{V}}q(v^{t-1}_{i}\mid v_{i}=v,v^{t}_{i})\;{\hat{p}}_{i}(v)$ (9)
-$p_{\theta}(e^{t-1}_{i,j}\mid G^{t},S)=\sum_{e\in\mathcal{E}}q(e^{t-1}_{i,j}\mid e_{i,j}=e,e^{t}_{i,j})\;{\hat{p}}_{i,j}(e)$
-Now sampling can be performed as Algorithm 2.
-It begins with graph $G^{T}$, where each node and edge is masked.
-At each step $t$, it calculates the distribution and samples the graph at step $t-1$, gradually reducing the noise and eventually resulting in a clean graph $G^{0}$.
-Our model can implicitly sample the molecular length through the empty token, which is crucial as target molecule may be unique for instructions.
-Training Objective
-Since the denoising model takes a noisy graph $G^{t}$ as input and aims to predict the clean graph $G^{0}$, we optimize the cross-entropy loss (represented as $\mathrm{CE}$) for each node and edge between the predicted probabilities $\hat{p}(G)$ and the clean graph $G$ during training.
-The diffusion loss is defined as:
-loss $=l(\hat{p}(V),V)+l(\hat{p}(E),E)$ (10)
-$=\sum_{1\leq i\leq m}\mathrm{CE}(v_{i},\hat{p}_{i}(V))+\sum_{1\leq i,j\leq m}\mathrm{CE}(e_{i,j},\hat{p}_{i,j}(E))$
-In addition, as previously discussed, maintaining balanced training across different modalities is crucial to prevent the model from diminishing its text comprehension ability after extensive training, as the tokens of the graph and text are non-overlapping.
-To address this issue, we incorporate a textual loss component to preserve the model’s text comprehension capability, which is defined as the cross-entropy loss for each token between the predicted logits $\hat{p}(S)$ and the ground-truth token in the instruction.
-Through the denoising model, the logits can be directly obtained alongside $\hat{p}(V)$ following a masked language model head, expressed as $\phi_{\theta}(G^{t},S)=[\hat{p}(S),\hat{p}(V),\hat{p}(E)]$.
-The final loss is thus represented as:
-$\text{loss}=l(\hat{p}(V),V)+l(\hat{p}(E),E)+l(\hat{p}(S),S)$ (11)
-See Algorithm 1 for the complete training procedure.
-Algorithm 1 Training Algorithm
-0:  A graph $G=(V,E)$ with description $S$
-1:  repeat
-2:     Sample $t\sim U(1,\dots,T)$
-3:     Sample $G^{t}\sim V\bar{Q}_{V}^{t}\times E\bar{Q}_{E}^{t}$
-4:     $\hat{p}(S),\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
-5:     loss = $l_{CE}(\hat{p}(E),E)+l_{CE}(\hat{p}(V),V)+l_{CE}(\hat{p}(S),S)$
-6:     optimizer.step(loss)
-7:  until converged
-Algorithm 2 Sampling Algorithm
-1:  $G^{T}$ with all masked nodes and edges
-2:  for $t=T$ to $1$ do
-3:     $\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
-4:     $p_{\theta}(v^{t-1}_{i}\mid G^{t},S)=\sum_{v}q(v^{t-1}_{i}\mid v_{i}=v,v^{t}_{i})\;{\hat{p}}_{i}(v)$
-5:     $p_{\theta}(e^{t-1}_{ij}\mid G^{t},S)=\sum_{e}q(e^{t-1}_{ij}\mid e_{ij}=e,e^{t}_{ij})\;{\hat{p}}_{ij}(e)$
-6:     $G^{t-1}\sim\prod_{i}p_{\theta}(v^{t-1}_{i}\mid G^{t},S)\prod_{ij}p_{\theta}(e^{t-1}_{ij}\mid G^{t},S)$
-7:  end for
-8:  return  $G^{0}$
+Under the cumulative transition matrices formulation $\bar{Q}_{V}^{t}=Q_{V}^{1}\dots Q_{V}^{T}$ and $\bar{Q}_{E}^{t}=Q_{E}^{1}\dots Q_{E}^{T}$, the forward process can also be expressed as $q(G^{t}\mid G^{0})=(V\bar{Q}_{V}^{t},E\bar{Q}_{E}^{t})$.
+With specific transition matrices, the 1-step and t-step marginals are in the closed form, and the posterior at time step t1t-1 admits a closed-form expression:
+$\begin{split}q(v_{i}^{t-1}|v_{i},v_{i}^{t})=(v_{i}^{t-1}{Q_{V}^{t}}^{\top}\circ v_{i}\bar{Q}_{V}^{t-1})/(v_{i}\bar{Q}^{t}_{V}{v_{i}^{t}}^{\top})\\ q(e_{ij}^{t-1}|e_{ij},e_{ij}^{t})=(e_{ij}^{t-1}{Q_{V}^{t}}^{\top}\circ e_{ij}\bar{Q}_{V}^{t-1})/(e_{ij}\bar{Q}^{t}_{V}{e_{ij}^{t}}^{\top})\end{split}$ (2)
+Graph Diffusion with [MASK] absorbing states
+Then we should focus on the formulation of noise model.
+Traditional methods consider the noise model using the marginal distributions over atom and bond types [47].
+While they claimed this makes training easier theoretically and experimentally, such assumptions are tailored to single-modal graph denoising model trained from scratch.
+In contrast, our goal is to integrate language understanding and graph denoising within a unified language-pretrained transformer, rendering this noise modeling assumption incompatible with our architecture.
+Therefore, we employ [MASK] absorbing states originally proposed in text diffusion [2] into our graph diffusion framework.
+Under this formulation, each element $v_{i}^{t}$ and $e_{i,j}^{t}$ independently decays into a [MASK] token as an absorbing state according to a noise schedule.
+Following the schedule $\beta(t)=(T-t+1)^{-1}$, the transition matrix $Q^{t}$ is formally expressed as follows for both nodes and edges simultaneously, where $z$ denotes the [MASK] index, and $x,y$ denote categories at consecutive timesteps:
+$[Q^{t}]_{x,y}$ $=\left\{\begin{array}[]{l l l}{{1}}&{{\text{if }}}{{x=y=z}}\\ {{1-\beta(t)}}&{{\text{if }}}{{x=y\not=z}}\\ {{\beta(t)}}&{{\text{if }}}{{y=z,x\not=z}}\\ {{0}}&{{\text{otherwise}}}\end{array}\right$.
+(7)
+The feasibility of this method is supported by the interpretation of BERT as a one-step discrete diffusion model [2], where the network is trained to predict the original tokens from [MASK] inputs.
+As we adopt a language-pretrained transformer familiar with this objective, the formulation aligns with its pretraining goal, thereby facilitating faster convergence of graph diffusion during fine-tuning.
+It also promotes better integration of text and graph modalities, and helps retain language understanding through direct parameter reuse.
+Empirical results further validate its effectiveness.
+3.4 Reverse Process and Denoising Network
+With the initial noisy graph $G^{T}$, the reverse process generates $G^{0}$ iteratively in reversed steps t=T,T1,,0t=T,T-1,...,0, formally represented as $p_{\theta}(G^{0:T-1}\mid G^{T},S)=\prod^{T}_{t=1}p_{\theta}(G^{t-1}\mid G^{t},S)$.
+It relies on a denoising neural network $\phi_{\theta}$ parameterized by $\theta$, which is trained to invert the forward process by predicting $G^{0}$ from $G^{t}$ conditioned on the instruction $S$ for each $t$, as the product over nodes and edges: $p_{\theta}(G^{0}\mid G^{t},S)=\prod_{v\in V}p_{\theta}(v\mid G^{t},S)\prod_{e\in E}p_{\theta}(e\mid G^{t},S)$.
+Then, to calculate the reverse distribution $p_{\theta}(G^{t-1}\mid G^{t},S)$, we adopt $x_{0}$-parameterization with noise back to the clean graph prediction, represented as $p_{\theta}(G^{t-1}\mid G^{t},S)=q(G^{t-1}\mid G^{t},\phi_{\theta}(G^{t},S))$, and marginalized over predictions of node and edge types $\mathcal{V^{\prime}},\mathcal{E^{\prime}}$:
+$\begin{split}&\!p_{\theta}(v^{t-1}_{i}\!\mid\!G^{t}\!,\!S)\!=\!\sum_{v\in\mathcal{V^{\prime}}}q(v^{t-1}_{i}\!\mid\!v_{i}\!=\!v,\!G^{t})\;p_{\theta}(v_{i}=v|G^{t},S)\\ &\!p_{\theta}(e^{t-1}_{i,j}\!\mid\!G^{t}\!,\!S)\!=\!\sum_{e\in\mathcal{E^{\prime}}}q(e^{t-1}_{i,j}\!\mid\!e_{i,j}\!=\!e,\!G^{t})\;p_{\theta}(e_{i,j}=e|G^{t},S)\end{split}$ (8)
+To enable graph tokens to selectively attend to text tokens, and allow textual representations to be refined by graph structure, we propose the Unified Text-Graph Transformer as our denoising network, which unifies two modalities within a single language-pretrained transformer.
+Below, we describe two key components of the model.
+Molecule Tokenization via Shared Vocabulary
+To adapt the language transformer for graph data, we first address the representation of molecular nodes.
+The key novel design is to treat molecular graph nodes as discrete tokens drawn from a shared vocabulary space, enabling seamless integration with transformer-based language models.
+Specifically, we assign a unique textual token (e.g., [C], [O-]) to each atomic category and embed them in the same token space.
+The initial embeddings are defined as $h^{0}=\text{Emb}(S+V)=[h_{1},\dots,h_{n},h_{n+1},\dots,h_{n+m}]\in\mathbb{R}^{(n+m)\times d_{h}}$, corresponding to $n$ instruction tokens and $m$ graph nodes.
+The graph node embeddings $h_{n+1},\dots,h_{n+m}$ are derived from the atom tokens (e.g., [[S], [O], [O]] for sulfur oxide).
+This tokenization strategy allows graph nodes to be processed identically to language tokens within a unified transformer, enabling attention-based cross-modal integration and consistent parameter sharing.
+It also prevents compositional ambiguities, such as mistakenly interpreting cobalt (Co) as a combination of carbon and oxygen.
 Type Model #params MACCS FTS $\uparrow$ RDK FTS $\uparrow$ Morgan FTS $\uparrow$ FCD $\downarrow$ Exact $\uparrow$ Valid $\uparrow$
-Specialist Auto-regressive (w.o. pretrain) T5-small 77M 0.704 0.578 0.525 2.89 0.064 0.608
-T5-base 248M 0.731 0.605 0.545 2.48 0.069 0.660
+Text Auto-regressive (w.o. pretrain) T5-base 248M 0.731 0.605 0.545 2.48 0.069 0.660
 T5-large 783M 0.823 0.731 0.670 1.22 0.279 0.902
 BioT5-base (reproduce) 252M 0.821 0.708 0.633 1.67 0.071 1.000
-Specialist Auto-regressive (pretrain) MolT5-small 77M 0.703 0.568 0.517 2.49 0.079 0.721
-MolT5-base 248M 0.721 0.588 0.529 2.18 0.081 0.772
+Text Auto-regressive (pretrain) MolT5-base 248M 0.721 0.588 0.529 2.18 0.081 0.772
 MolT5-large 783M 0.834 0.746 0.684 1.20 0.311 0.905
@@
 tgm-dlm w/o corr 125M 0.874 0.771 0.722 0.89 0.242 0.789
-LLM chatgpt3.5 (0-shot) 7B 0.703 0.568 0.517 2.49 0.079 0.721
-Chatgpt3.5 (10-shot) 7B 0.847 0.708 0.624 0.57 0.139 0.887
+LLM Chatgpt3.5 (0-shot) - 0.703 0.568 0.517 2.49 0.079 0.721
+Chatgpt3.5 (10-shot) - 0.847 0.708 0.624 0.57 0.139 0.887
+gpt4 (0 shot) - 0.787 0.577 0.492 1.33 0.055 0.883
+gpt4 (10 shot) - 0.872 0.736 0.659 0.66 0.092 0.891
+LlaSMol-Mistral 7B 0.853 0.726 0.650 0.70 0.132 0.935
 Graph-based Momu-S 113M 0.244 0.103 0.047 22.21 0.000 1.000
+Digress (similarity guidance) 289M 0.577 0.389 0.288 28.82 0.014 0.854
+Graph-DiT 162M 0.374 0.269 0.159 18.58 0.000 0.909
+3M-Diffusion 162M 0.548 0.370 0.273 4.42 0.005 1.000
 UTGDiff (w.o. pretrain) 125M 0.867 0.763 0.695 0.92 0.227 0.856
 UTGDiff (pretrain) 125M 0.885 0.795 0.724 0.86 0.374 0.893
-Table 1: Results of the instruction-based molecule generation on ChEBI-20 dataset for both with/without pretraining setting.
-Pretraining Method
-Recent studies have demonstrated that BERT can be viewed as a one-step [MASK] absorbing diffusion model  (Devlin et al.
-2019; Austin et al.
-2021), since one-step noising replaces some tokens with [MASK] and denoising performs masked prediction.
-This shared training objective allows for the effective initialization from pre-trained language models (He et al.
-2022).
-Therefore, to preserve language comprehension, we utilize the pre-trained RoBERTa model (Liu et al.
-2019) for initialization.
-However, the introduction of novel graph tokens during tokenization renders the direct loading of pre-trained BERT/RoBERTa parameters insufficient for mask prediction on graph tokens.
-Consequently, our pretraining methods are designed to equip the model with mask prediction capabilities for graph data while retaining its functionality for text data.
-We collect paired or single-modal data in both text and molecular graph modalities, mask all tokens and edge indices with a 15% probability, and pretrain the model using the masked language prediction method.
-The effectiveness of this pretraining approach for graph data has been validated by previous studies on graph pretraining (Hu et al.
-2020; Hou et al.
-2022).
-Experiment
-Dataset
+Table 1: Results of the instruction-based molecule generation on ChEBI-20 for both with/without pretraining setting.
+Attention Bias
+Next, we address the representation of edges.
+In the attention layer with parameters for values, keys, queries, and outputs, i.e., $W_{V},W_{Q},W_{K}\in\mathbb{R}^{d_{h}\times d_{k}},W_{O}\in\mathbb{R}^{d_{k}\times d_{h}}$, we introduce an additional bias term to incorporate edge information.
+The attention score between the $i$-th and $j$-th graph tokens is now formalized as:
+$\begin{split}\hat{A}_{i,j}^{l}=\frac{1}{\sqrt{d_{k}}}\left(h_{i}^{l}W_{Q}\right)\left(h_{j}^{l}W_{K}\right)^{\top}+b_{i,j}^{l},\\ A^{l}=\mathrm{softmax}(\hat{A}^{l}),\ \text{Attn}(H)=A^{l}H^{l}W_{V}W_{O}\end{split}$ (9)
+Here, $b_{i,j}$ denotes the bias added to the attention scores between graph tokens to incorporate structural information [54, 57].
+To ensure the expressive power in distinguishing neighbors as one-layer GNNs and the symmetry between edge representations, we therefore represent the initial embedding from the edge category as $b_{i,j}^{0}$ and the update of attention bias as follows:
+[…]
+The FFN module applied to attention scores follows the structure in standard transformer, including a residual connection and LayerNorm.
+Our design distinguishes neighbors using edge-specific initial biases $b_{i,j}^{0}$, and enforces symmetry by symmetrizing the attention scores.
+Beyond structural constraints, our model minimally modifies standard transformer, enabling joint text-graph processing without additional GNNs and allowing full reuse of pre-trained language model weights.
+Also, layer-wise text-graph interactions enable graph tokens to selectively attend to text tokens and text representations to be refined by graph, supporting fine-grained interactions.
+Output
+After obtaining the final hidden states $h^{L}$ for tokens and $b^{L}$ for edges through stacked attention layers, we map them back to the vocabulary space for the probability of each category through two separate masked language model heads at the end of the transformer.
+This yields the node logits $\log p_{\theta}(v_{i}\mid G^{t},S)$ and edge logits $\log p_{\theta}(e_{i,j}\mid G^{t},S)$ for the clean graph distribution.
+Assembling them into matrices $\hat{p}(V)\in\mathbb{R}^{m\times a}$ and $\hat{p}(E)\in\mathbb{R}^{m\times m\times b}$, the unified denoising network is summarized as:
+$\phi_{\theta}(G^{t},S)=\hat{p}(G)=[\hat{p}(V),\hat{p}(E)]$ (11)
+Sampling
+Empirically, positional embeddings can improve performance in the unified transformer by atom locations, while inherently break permutation equivariance.
+Therefore, inspired by permutation-based graph encoding methods [37, 39], we adopt a sampling strategy that preserves equivariance at inference: positional embeddings are applied to a fixed node order during training, while random permutations of position indices are introduced during sampling.
+Therefore, for any permutation $\pi$, the model generates a graph with node $V$, adjacency matrix $A$, and position indices $i$ with equal probability, satisfying $\mathbb{P}_{i}(V,A)=\mathbb{P}_{\pi^{T}i}(\pi^{T}V,\pi^{T}A\pi)$, thereby ensuring equal probability over the molecule’s permutation set during sampling.
+We prove the equivariance in Appendix B in the supplementary material [52].
+3.5 Training Objective
+To train the denoising network, the original method [47] optimizes the cross-entropy loss (denoted as $\mathrm{CE}$) between the predicted probabilities $\hat{p}(G)$ and the clean graph $G$ in the context of single-modal graph diffusion.
+However, in our multimodal setting, such graph-only supervision is insufficient.
+Since the model is required to interpret textual instructions while generating molecular structures, an additional loss term related to the text modality is necessary to preserve its language understanding capabilities.
+Therefore, we introduce a novel additional textual loss by applying the cross-entropy objective between the predicted logits $\hat{p}(S)$ and the ground-truth instruction tokens.
+Through the denoising transformer, the logits $\hat{p}(S)$ can obtained alongside $\hat{p}(V)$ and $\hat{p}(E)$ via the masked language model head, formulated as $\phi_{\theta}(G^{t},S)=[\hat{p}(S),\hat{p}(V),\hat{p}(E)]$.
+Then we define the final training loss $\mathcal{L}$ as:
+$\begin{split}\mathcal{L}&=l(\hat{p}(V),V)+l(\hat{p}(E),E)+l(\hat{p}(S),S)\\ \!&=\!\sum_{1\leq i\leq m}\!\mathrm{CE}(v_{i},\!\hat{p}_{i}(V))\!+\!\sum_{1\leq i,j\leq m}\!\mathrm{CE}(e_{i,j},\!\hat{p}_{i,j}(E))\\ &+\!\sum_{1\leq i\leq n}\!\mathrm{CE}(S_{i},\!\hat{p}_{i}(S))\end{split}$ (12)
+Model Validity $\uparrow$ Distribution Learning Condition Control
+(w/o rule checking) Coverage $\uparrow$ Diversity $\uparrow$ Similarity $\uparrow$ Distance $\downarrow$ Synth.
+$\downarrow$ O2Perm $\downarrow$ N2Perm $\downarrow$ CO2Perm $\downarrow$ Avg.
+MAE $\downarrow$
+JTVAE-BO 1.0000 (N.A.) 10/11 0.7366 0.7294 23.5990 1.0714 1.0781 1.2352 1.0978 1.1206
+DiGress 0.9913 (0.2362) 11/11 0.9099 0.2724 22.7237 2.9842 1.7163 2.0630 1.6738 2.1093
+GDSS 0.9205 (0.9076) 9/11 0.7510 0.0000 34.2627 1.3701 1.0271 1.0820 1.0683 1.1369
+Graph-DiT-LCC 0.9753 (0.8437) 11/11 0.8875 0.9560 7.0949 1.3099 0.8001 0.9562 0.8125 0.9697
+Graph-DiT 0.8245 (0.8437) 11/11 0.8712 0.9600 6.6443 1.2973 0.7440 0.8857 0.7550 0.9205
+UTGDiff (Ours) 0.9596 (0.9085) 11/11 0.8532 0.9637 6.0923 1.0208 0.6462 0.7826 0.6459 0.7739
+Table 2: Multi-Conditional Generation of 10K Polymers: Results on the synthetic score (Synth.) and three numerical properties (gas permeability for O2, N2, CO2).
+MAE is calculated between the input conditions and the properties of the generated polymers using Oracles.
+Best results are highlighted.
+This loss helps preserves the model’s language comprehension by reinforcing masked language modeling behavior.
+When a text token is provided as input, the model is explicitly trained to predict the same token, ensuring token-level consistency.
+This aligns with the core objective of MLM: reconstructing masked tokens from context demonstrates semantic comprehension, while preserving the identities of unmasked tokens stabilizes the language representations and supports the retention of language understanding during training.
+3.6 Pretraining Method
+We have mentioned that BERT can be interpreted as a one-step [MASK] absorbing diffusion model which discussed in [2].
+This shared training objective enables initialization from pre-trained language models [18].
+Accordingly, we initialize our model from the pre-trained RoBERTa [34] to preserve language comprehension.
+However, RoBERTa initialization is insufficient for masked prediction on novel graph tokens.
+To address this, we further pretrain the model on masked graph data while preserving its text modeling ability.
+We collect paired and single-modal data from textual and graph modalities, randomly mask 15% of tokens and edge indices, and train the model with masked language modeling.
+The effectiveness of this method on graph data has been validated by prior work [20].
 Model names
@@
 UTGDiff Pubmed, ZINC, pubchem324K
-Table 2: Review of pretraining datasets. Some data listed in BioT5 is unavailable, and we reproduce it with our dataset.
-Pretraining Dataset
-We collect several open-source datasets for pretraining.
-For the molecule-text pair data, we select the PubChem-324K dataset (Liu et al.
-2023c), which contains approximately 320,000 molecule-text pairs from PubChem.
-For the single-modal data, we utilize the molecular graph dataset ZINC20 (Irwin et al.
-2020) and biomedical text data from PubMed abstracts (White 2020), totaling nearly 100 million entries.
-Due to computational budget constraints, we are unable to scale up the training size and batch size of the pretraining corpus to the extent achieved by BioT5 (Pei et al.
-2023), which utilizes non-open-source pretraining data.
-Generation Dataset
-For molecule generation tasks, we utilize the ChEBI-20 dataset (Edwards, Zhai, and Ji 2021) for finetuning and evaluation, which consists of 33,010 entries, with 10% allocated for validation and 10% for testing.
-We adhere to the original data split settings.
-Editing Dataset
-For molecule editing, we select the retrosynthesis and forward reaction prediction tasks from Mol-Instructions (Fang et al.
-2023).
-The retrosynthesis task begins with a target compound and identifies potential precursor molecules as outputs.
-The forward reaction prediction task involves forecasting likely products as output molecular graphs, given specific reactants and reagents.
-Each of these two datasets contains approximately 120,000 reactant-product pairs for training and 1,000 pairs for testing, with each entry accompanied by an instruction.
-Figure 3: Some generation examples from our model and baseline models.
-model 1-64 65-96 97-128 129-160 $>$ 160
-BioT5 (reproduce) 0.810 0.866 0.859 0.817 0.788
+Table 3: Review of pretraining datasets.
+Specifically, for multimodal data, we use the PubChem-324K [35], containing 320K molecule-text pairs.
+For single-modal data, we employ molecular graphs from ZINC20 [22] and biomedical text from PubMed abstracts [50], totaling nearly 100 million entries.
+Since PubChem-324K is explicitly curated to avoid overlap in generation and others are single-modal data, there’re no risk of data leakage.
+4 Experiment
+4.1 Instruction-based Molecule generation
+First, we evaluate UTGDiff on molecule generation tasks, to demonstrate whether UTGDiff achieves high instruction fidelity and structural validity under structural descriptions.
+Dataset
+We utilize the ChEBI-20 dataset [11] for fine-tuning and evaluation. It contains 33,010 molecule-instruction pairs, with 10% allocated for validation and 10% for testing.
+Baselines
+Baseline models include (1) Sequence-based models—T5, MolT5 [12] and BioT5 [41]; (2) Large language models—ChatGPT-3.5, GPT-4 (2024-04-09), and LlaSMol-Mistral-7B fine-tuned on molecular data; (3) text diffusion model tgm-dlm [15]; and (4) graph-based models—MoMu-S [46] (flow-based graph decoder), Digress[47] (predictor-guided diffusion), Graph-DiT [31] (predictor-free multi-condition diffusion), and 3M-Diffusion [59] (separate encoders/decoders).
+BioT5 is retrained on our pretraining data for fair comparison (Table 3) due to unavailable original pretraining data.
+Evaluation
+We adopt 6 metrics to assess the similarity and quality of generated molecules: (1) Exact match (Exact); (2) Molecular validity (Validity); (3) 3 fingerprint metrics for similarity in functional group level (FTS); (4) Fréchet ChemNet Distance for similarity in set level (FCD).
+The full description is in Appendix C in the supplementary material [52].
+Results
+The results are summarized in Table 1.
+Additional results with different random seeds and hyperparameter details are provided in Appendix C and D in the supplementary material [52].
+Our main findings are as follows:
+UTGDiff outperforms all text-based autoregressive methods on similarity metrics with comparable pretraining data size.
+It achieves substantial improvements in Exact match and FTS scores, surpassing text baselines by 6% on average, highlighting the diffusion paradigm’s strength in capturing complex molecular structures.
+UTGDiff significantly outperforms graph-based methods.
+Previous graph diffusion approaches struggle with complex conditional generation due to information bottleneck and limited scalability.
+In contrast, UTGDiff enables finer-grained integration of language and graph modalities through unified model and layer-wise aggregration, leading to substantial improvements in instruction-following ability.
+Simple text diffusion methods fall short without deep text-graph integration.
+UTGDiff improves the Exact match score over 15% compared to text diffusion models, indicating that our gains are not solely attributable to the diffusion framework itself but to the deep structural integration of graphs within the language model.
+UTGDiff achieves better performance with significantly fewer parameters. Our model contains only 125M parameters, substantially smaller than other models but with better performance.
+UTGDiff is marginally outperformed by BioT5+ and a few fine-tuned LLM [29], which can be attributed to their much larger pretraining corpora.
+Minor shortcomings in molecular validity stem from inevitable valence violations, while lower FCD can due to the metric’s inability to distinguish cross-instruction mismatches, with further analysis in Appendix A in the supplementary material [52].
+Figure 3: Some generation examples comparison on CHEBI-20.
+Case Study
+Several examples are shown in Figure 3.
+UTGDiff successfully generates accurate molecules as baseline models (Example 1), and can produce correct structures where other models fail (Example 2).
+Furthermore, it can capture critical instruction details more effective, such as "carboxylic acid group of glycine", demonstrating its ability for contextually appropriate outputs (Example 3).
+Length 1-64 65-96 97-128 129-160 >> 160
+BioT5 0.810 0.866 0.859 0.817 0.788
 UTGDiff 0.860 0.903 0.901 0.853 0.821
-Table 3: Analysis for scalability on instruction length.
+BertzCT <100 100-300 300-800 >800
+BioT5 0.765 0.863 0.831 0.857
+tgl-dlm 0.757 0.859 0.841 0.882
+UTGDiff 0.837 0.898 0.881 0.889
+Table 4: Analysis for scalability of the model
+Scalability
+We further assess UTGDiff’s scalability on instruction and molecule complexities.
+We report MACCS FTS across instruction length and BertzCT intervals, a topological metric reflecting graph size and branching.
+Table 5 shows that UTGDiff maintains strong performance across all complexity levels, demonstrating robustness on detailed instructions and complex molecules.
+Model MACCS $\uparrow$ RDK $\uparrow$ Morgan $\uparrow$ FCD $\downarrow$ Exact $\uparrow$ Valid $\uparrow$
+MolT5-base 0.733 0.651 0.621 2.20 0.204 0.761
+BioT5-base 0.737 0.646 0.595 1.40 0.242 1.000
+tgm-dlm 0.741 0.667 0.612 2.06 0.221 0.765
+3M-Diffusion 0.495 0.332 0.242 3.47 0.015 1.000
+UTGDiff 0.763 0.675 0.623 2.03 0.386 0.847
+Table 5: Analysis for generalization in zero-shot experiment
+Generalization
+To assess the model’s generalization to unseen instructions, we conduct a zero-shot evaluation on PCDes using the model trained only on ChEBI.
+Sentence-BERT confirms that PCDes instructions differ from ChEBI’s, with 0.929 cosine similarity compared to 0.999.
+UTGDiff performs competitively across similarity metrics and generalizes well without retraining.
+4.2 Multi-Property Conditional Molecule Generation
+The instruction dataset above mainly focuses on molecular structures.
+However, real-world molecular generation scenarios often require inverse design, based on specified property conditions.
+To better reflect this, we evaluate multi-property conditional generation, formulating numerical properties as part of instructions to assess our model’s ability across more practical conditions.
+We adapt UTGDiff by expressing property constraints textually and introducing separate MLP encoders to map each condition into an input embedding.
+Dataset
+We utilize the polymer dataset for materials, which includes three numerical gas permeability properties: O2Perm, CO2Perm, and N2Perm.
+It contains 553 entries, with 20% for validation and 20% for testing following the data split protocol in [31].
 Baselines
-Several prior studies have explored sequence-based models for evaluation.
-For the ChEBI-20 dataset, the baseline models include (1) specialist auto-regressive models such as T5 (Raffel et al.
-2020), MolT5 (Edwards et al.
-2022) and BioT5 (Pei et al.
-2023); (2) LLMs such as ChatGPT-3.5 for zero-shot and 10-shot; and (3) text diffusion model tgm-dlm (Gong et al.
-2024).
-For the editing datasets, the baseline models include (1) specialist auto-regressive models such as BioT5 (Pei et al.
-2024) and TEXT+CHEM T5 (Christofidellis et al.
-2023); (2) LLM such as InstructMol (Cao et al.
-2023) and other models.
-Since BioT5’s pretraining data is not open-sourced, we reproduce their experiments training on our pretraining datasets using reported hyperparameters, rather than referencing their reported performance.
-These baselines help us highlight the deficiencies of sequence-based models and clarify our improvements.
-Additionally, to assess the effectiveness of our diffusion method, we finetuned MoMu-S (Su et al.
-2022) on the CHEBI-20 dataset for instruction-based generation.
-Due to the lack of capability to interpret language in the graph-only structure-aware network, there’re few baselines related to graph diffusion and it’s challenging to directly adapt them into these tasks.
-Therefore, we don’t consider them as baselines, but further discuss these methods in Ablation Study.
-Type Model MACCS RDK Morgan Exact Valid
-LLM LLAMA 0.029 0.018 0.017 0.000 0.010
-GALACTICA 0.274 0.167 0.134 0.000 0.986
+We compare against several baselines, including the VAE-based JTVAE model with Bayesian optimization (JTVAE-BO), and diffusion-based methods such as predictor-guided GDSS, DiGress [47], and predictor-free Graph-DiT [31].
+We follow the baseline setup in Graph-DiT [31], restricting comparisons to graph-based models, as text-based models are not designed for precise numerical control on property-conditioned generation.
+Evaluation
+For fair comparison, we adopt the same evaluation protocol on 10,000 generated molecules.
+In addition to validity and FCD metrics described above, we have: (1) heavy atom type coverage (Coverage); (2) diversity among examples (Diversity); (3) fragment-based similarity with the reference set (Similarity); (4) MAE between the generated and conditioned molecules, including synthetic accessibility score (Synth.) and MAE for numerical conditions (Property).
+Results
+Our findings are shown in Table 2 and described as follow:
+UTGDiff maintains strong molecular distribution modeling beyond prior diffusion methods.
+UTGDiff achieves the best performance in both similarity and distribution-level metrics.
+This is notable since it contains only 11 heavy atom types and 5 condition types, much simpler than the diversity seen during pretraining.
+These results highlight UTGDiff’s generalization from large-scale structure-focused pretraining to property-conditioned generation.
+UTGDiff advances conditional graph diffusion and achieves SOTA multi-property control.
+Across all three properties, UTGDiff consistently reduces MAE, averaging a 19% improvement over the previous SOTA Graph-DiT.
+This demonstrates that our layer-wise interaction and dedicated numerical encoding enable better multi-property control, validating UTGDiff’s scalability for more diverse property conditions and reflecting real-world generation scenarios.
+Figure 4: Some generation examples comparison on O2-N2-CO2 Perm.
+Case Study
+Several examples are shown in Figure 4.
+UTGDiff successfully generates near-identical molecules that match the ground truth and outperform baseline models (Example 1).
+Moreover, it accurately recovers correct atom types (S,Si) under property constraints where other models fail (Example 2 and 3), highlighting UTGDiff’s precise understanding of numerical conditions.
+4.3 Instruction-Based Molecule Editing
+We further evaluate UTGDiff on molecule editing tasks to assess generalization in diverse scenarios.
+Unlike generation, editing requires an additional source graph as input, incorporated into the instruction in same graph format as output.
+Evaluation follows the same metrics as in instruction-based molecule generation.
+Dataset
+We select the retrosynthesis and forward reaction prediction tasks from Mol-Instructions [13].
+The retrosynthesis task predicts precursors from a target compound, while the forward reaction prediction task forecasts products from given reactants and reagents.
+Each dataset contains 120K training pairs and 1,000 testing pairs.
+Baseline
+For the editing datasets, the baseline models include (1) specialist auto-regressive models such as BioT5 [41] and TEXT+CHEM T5 [5]; (2) LLM such as InstructMol [4] and others.
+We also reproduce BioT5’s experiments on our pretraining datasets rather than referencing their reports as discussed above.
+Results
+The results are presented in Table 6 and Table 7.
+[…]
+LLM GALACTICA 0.274 0.167 0.134 0.000 0.986
 Mol-Instructions 0.487 0.283 0.230 0.009 1.000
@@
 T5 TEXT+CHEM T5 0.765 0.685 0.585 0.141 0.698
-BioT5+ (reproduce) 0.904 0.843 0.810 0.480 1.000
+BioT5 (reproduce) 0.904 0.843 0.810 0.480 1.000
 UTGDiff 0.904 0.847 0.817 0.541 0.945
 Table 6: Mol-Instruction Retrosynthesis results.
-Instruction-Based Molecule Generation
-Results
-Our results are shown in Table 1.
-The evaluation metrics are chosen based on previous studies, including a validity metric and five graph-level similarity metrics, with the exception of text-level similarity metrics such as BLEU.
-Here, we present the results of a single experiment.
-Detailed descriptions of these metrics, additional results and variance obtained from different seeds, and details for experiments such as hyperparams are available in the supplementary material C,D.
-Our experimental findings are as follows:
-Our model outperforms all auto-regressive methods across almost all similarity metrics with the same level of pretraining corpus.
-UTGDiff exhibits substantial improvements in the Exact match score, with gains of at least 6% over the best auto-regressive baseline.
-Also, UTGDiff achieves the highest FTS score, surpassing baseline models by at least 2%.
-These results indicate the methodological advantages of diffusion paradigm, since it is better suited to capturing the complex structures inherent in molecules.
-Compared to text diffusion models, UTGDiff also performs better in both similarity and validity metrics.
-Especially, UTGDiff exhibits a remarkable 15% improvement in the Exact match score.
-This indicates that the superior performance of our model is not solely attributable to the diffusion paradigm.
-Instead, by incorporating graph-based representations, our model leverages explicit structural information in graph data processing, which is not attainable with text diffusion methods.
-Type Model MACCS RDK Morgan Exact Valid
-LLM LLAMA 0.002 0.001 0.001 0.000 0.010
-GALACTICA 0.127 0.036 0.051 0.000 0.995
+[…]
+LLM GALACTICA 0.127 0.036 0.051 0.000 0.995
 Mol-Instructions 0.509 0.313 0.262 0.045 1.000
@@
 T5 TEXT+CHEM T5 0.789 0.705 0.652 0.141 0.698
-BioT5+ (reproduce) 0.954 0.907 0.890 0.684 1.000
+BioT5 (reproduce) 0.954 0.907 0.890 0.684 1.000
 UTGDiff 0.973 0.943 0.942 0.825 0.972
 Table 7: Mol-Instruction forward reaction prediction results.
-Our model is far better than the graph-based MoMu-S model.
-The low performance of MoMu-S may be attributed to the limitations of MoFlow, as it struggles to follow complex instructions and generate long molecules, showing the importance of unified diffusion.
-Our model has fewer parameters than other models with comparable performance. Our model has only 125M parameter size, smaller than other T5 models and LLMs.
-Our model performs marginally worse than the reported results of BioT5+ and the recently finetuned LLM (Li et al.
-2024).
-This can be attributed to the larger pretraining corpora they used as discussed in the ablation study.
-Also, our model shows minor shortcomings in validity metrics due to inevitable valence-induced invalidity.
-The low value in FCD can be attributed to misleading influences from other generated molecules, as discussed in supplementary material A.
-Case Study
-We present several generated examples in Figure 3.
-The model successfully generates accurate results comparable to baseline models (Example 1).
-In some instances, it produces the correct solution when other models fail (Example 2).
-Also, it can effectively capture essential prompt words like ”tosylate” and generate the required sulfur atom, demonstrating its capability to produce contextually appropriate outputs (Example 3).
-Scalability on Instruction
-We also present an experiment assessing performance with varying complexities of instructions.
-We report the MACCS results for varying lengths of instructions and compare them with BioT5.
-The results indicate that our diffusion model achieves better outcomes across all intervals, demonstrating enhanced scalability when handling more complex instructions.
-Instruction-Based Molecule Editing
-We also conducted experiments on molecule editing tasks to demonstrate the model’s generalization capabilities across diverse tasks.
-Transferring from molecular generation to editing requires an additional source molecular graph as input, which is incorporated into the instruction input in graph format.
-The results are presented in Table 4 and Table 5.
+UTGDiff consistently outperforms all baselines across fingerprint similarity and exact match metrics, achieving 5% improvement in average in two tasks.
+These results underscore the methodological advantages of our diffusion approach over autoregressive baselines.
+Compared to InstructMol-G and InstructMol-GS, we achieves a twofold increase in exact match scores on forward reaction prediction, along with at least a 5% improvement across all metrics on both tasks.
+This highlights the effectiveness of graph diffusion generation, as InstructMol-G integrates a 2D graph encoder but still relies on a sequence-based decoder, limiting its structural modeling capacity.
+4.4 Ablation Study
+We ablate our model on the model design and training strategy on CHEBI-20 dataset, validating the contribution of these key designs.
+Model Design
+We conduct ablation studies focusing on two aspects of model design: (1) the contribution of the diffusion module to generation performance; and (2) the effectiveness of key architectural components, including unification via attention bias and the absorbing state implemented through [MASK] tokens.
 […]
@@
 1000 0.881 0.901 11.388
-BioT5 - 0.843 1.000 0.763
-Table 6: Ablation study on diffusion framework. Steps = 1 is equal to without diffusion
+Table 8: Ablation on diffusion. Steps = 1 means without diffusion.
+(1) Effect of the Diffusion Module.
+The $x_{0}$-parameterization allow us to use flexible step-size inference during generation.
+Performing generation with one-step logits represents a setting without diffusion.
+As shown in Table 8, the diffusion module improves performance, demonstrating the importance of progressive refinement.
 […]
-w.o. unifying 0.583 28.47 0.753 -
-3M-Diffusion - - - 0.871
-UTGDiff 0.867 0.923 0.856 0.957
-Table 7: Ablation study on unifying denoising network.
-Our model outperforms baseline models in all fingerprint and exact match metrics, with gains of at least 2% in forward reaction prediction tasks.
-These results underscore the methodological advantages of our model over previous string-based auto-regressive models.
-When compared to InstructMol-G and InstructMol-GS, our model exhibits a remarkable doubling of the exact match score in forward reaction prediction, with at least a 5% improvement in all metrics across both tasks.
-These results highlight the significance of the diffusion generation method for these tasks, as InstructMol-G incorporates a 2D molecular graph encoder while retaining a sequence-based decoder approach.
-Ablation Study
-In our ablation study, we validate our model focusing on the impact of pretraining and model design. All ablation experiments were performed using the CHEBI-20 dataset.
-Model Design
-Our investigation of model design focuses on two key components to highlight the importance of novel modules: (a) Ablation of the diffusion module.
-Using the $x_{0}$-parameterization, we can perform inference with flexible step size during the reverse process, while the one-step logits represent a scenario without diffusion.
-The effectiveness of the diffusion module is demonstrated in Table 6.
-(b) Ablation of the unified denoising network.
-This experiment involves the independent use of the graph diffusion network and text encoder, guided by a similarity gradient (Liu et al.
-2023b), without the unification of the network.
-Additionally, we draw from the recent 3M-Diffusion model (Zhu, Xiao, and Honavar 2024), a text-guided graph diffusion method that employs separate encoders, and evaluate our generation results using their Similarity metrics.
-The results, shown in Table 7, indicate the effectiveness of the unified network.
+Cross-attn 0.693 2.720 0.632
+marginal absorbing 0.877 1.852 0.837
+UTGDiff 0.885 0.866 0.893
+Table 9: Ablation on our model design.
+(2) Effect of Architecture.
+To ablate the unification strategy, we replace attention bias with cross-attention, encoding instructions separately for cross-modal fusion.
+To ablate the absorbing state design, we replace the [MASK]-based noise representation with marginal distributions as Graph-DiT [31].
+Table 9 confirms the effectiveness of both attention bias and [MASK]-based noise modeling.
 Model MACCS $\uparrow$ FCD $\downarrow$ Valid $\uparrow$ Train loss $\downarrow$ Entry
+All 0.885 0.866 0.893 7.69e-4 100M
+Pair-only 0.870 0.892 0.879 1.99e-3 300K
 from scratch 0.867 0.923 0.856 1.58e-2 0
-Pair-only 0.870 0.892 0.879 1.99e-3 300K
-All 0.885 0.866 0.893 7.69e-4 100M
+no textual loss 0.882 1.224 0.833 3.57e-3 100M
+no position index 0.764 2.387 0.821 2.92e-1 100M
 Table 10: Ablation study on pretrain data.
-Pretraining
-We conducted another ablation experiment to assess the impact of pretraining scale.
-The comparison, detailed in Table 8, highlights the critical roles of both single-modal and paired pretraining datasets and demonstrates the importance of scale, particularly concerning the validity.
-Conclusions and Limitations
-In this paper, we propose UTGDiff, an instruction-based framework that unifies graph and text data within a single transformer through attention bias, leveraging for the denoising generation of graph data.
-Our model is evaluated on tasks involving instruction-based molecule generation and editing, showing superior performance while requiring fewer parameters with the same level of pretraining corpus.
-Several limitations constrain this research.
-A critical constraint is our inability to scale the pretraining corpus to a larger size, a factor that significantly benefits models like BioT5 and may have limited our model’s ability to generalize across different molecular structures.
-Additionally, further experiments on more advanced discrete diffusion frameworks should be explored.
-Supplementary Material
-A. Analysis of FCD
+Training Strategy
+We conducted ablation experiments to assess the impact of pretraining scale and the textual loss $l(\hat{p}(S),S)$ during fine-tuning.
+Table 10 highlights the importance of both single-modal and paired pretraining datasets, demonstrating that scaling pretraining data significantly benefits downstream performance.
+The comparison of textual loss shows that incorporating textual loss leads to better convergence and higher validity.
+The comparison of position index shows the necessity of our sampling stratage.
+5 Conclusions and Limitations
+We propose UTGDiff, a unified instruction-based framework that integrates graph and text within a single transformer.
+With absorbing noise, atom-specific tokens, and attention bias, it enables fine-grained graph diffusion generation conditioned on complex instructions.
+Our model excels in molecule generation and editing, covering instructions for molecular structures or properties, while maintaining a smaller parameter size compared to existing baselines.
+Despite these advancements, UTGDiff is limited by the scale of its pretraining corpus, which may constrain generalization to more diverse molecular structures.
+Future work will focus on scaling pretraining and exploring more advanced discrete diffusion techniques to further enhance instruction-conditioned generation.
+Appendix A Analysis of FCD
 FCD is a metric originally used to measure the distance between two sets of distributions.
@@
 For instance, the instruction $S_{1}$ match the molecule ‘CCO’, and instruction $S_{2}$ match the molecule ‘[He]’.
-If we generate ‘CC’ with instruction $S_{1}$ and ‘[He]’ with instruction $S_{2}$, there’s no doubt that FCD([‘CCO’, ‘[He]’], [‘CC’, ‘[He]’]) ¿ 0.
+If we generate ‘CC’ with instruction $S_{1}$ and ‘[He]’ with instruction $S_{2}$, there’s no doubt that FCD ([‘CCO’, ‘[He]’], [‘CC’, ‘[He]’]) > 0.
 However, If we generate ‘[He]’ with instruction $S_{1}$ and ‘CCO’ with instruction $S_{2}$, which is a totally irrelavant results, we’ll got an unexpected result FCD ([‘CCO’, ‘[He]’], [‘[He]’, ‘CCO’]) = 0.
@@
 The difference in magnitude between the two settings is an inherent property of FCD, which is related to the number of molecule sets.
-B. Proof of required property
-The edge attribution must maintain symmetry.
-This is conducted since the start of embedding is symmetrical as they have same categories.
-And therefore the edge attribution maintains symmetry in each layer.
-Another required property for graph generation model is to ensure that the node’s attribution in the graph is equivariant to reorderings, meaning that the matrices under reordering can represent the same graph.
+Appendix B Proof of equivarient property
+The graph generation model needs to ensure that the node’s attribution in the graph is equivariant to reorderings, meaning that the matrices under reordering can represent the same graph.
 This property requires no positional embedding in the denoising network.
@@
 The theorem to be proven in this section can be written as follows:
-Theorem 1
+Theorem 1.
 (Equivariancy for Graph Generation) For any permutation $\pi$, the model will generate a graph with node feature $V$ and adjacency matrix $A$ with input position index $i$, satisfying $P_{i}(V,A)=P_{\pi^{T}i}(\pi^{T}V,\pi^{T}A\pi)$.
 The proof aims to show that: If the model’s network architecture is equivariant with input mapping permutation and the training loss is permutation invariant, the model will generate a distribution reordering with position index.
-Lemma 1
+Lemma 2.
 […]
-Proof 1
+Proof.
 Since the defined loss function is the same for each node and edge, we have
@@
 $=l(\hat{G},G)$
-Lemma 2
+Lemma 3.
 […]
-Proof 2
+Proof.
 Define $G_{t}=(X_{t},E_{t})$ as the noisy graph, $(\pi.X_{t},\pi.E_{t})$ as the permutation. Since the input is permutation equivariant, and we have:
@@
   • The Layer normalization is permutation equivariant.
-Therefore, the model is a combination of permutation equivariant blocks.
-Proof 3
+Therefore, the model is a combination of permutation equivariant blocks. ∎
+Proof.
 proof of theorem 1
 For the input embedding after model permutation, we have $\pi^{T}X+\pi^{T}emb(i)=\pi^{T}(X+emb(i))$, Therefore, the assumption of input mapping permutation equivariant holds.
-Following the proof in DiGress, with the lemma(Xu et al.
-2022): if a distribution $p(G_{T})$ is invariant to the action of a group G and the transition probabilities $p(G_{t-1}|G_{t})$ are equivariant, them $p(G_{0})$is invariant to the action of G.
+Following the proof in DiGress, with the lemma: if a distribution $p(G_{T})$ is invariant to the action of a group G and the transition probabilities $p(G_{t-1}|G_{t})$ are equivariant, them $p(G_{0})$is invariant to the action of G.
 We apply this result to the special case of permutations:
@@
 Table 11: FCD comparison
-C. Details for evaluation metrics
+Appendix C Details for evaluation metrics
 Since the model generates a matrix of nodes and edges, although ultimately a SMILES will be parsed based on the matrix, it is not directly generated, which can easily cause the problem of different strings for same graph.
-Therefore, although molecules can be represented by biological sequence structures, and previous models have also established evaluation indicators from the level of string similarity, this article does not involve these indicators in comparison, including NLP indicators such as BLEU(Papineni et al.
-2002), Levenshtein(Miller, Vandome, and McBrewster 2009).
+Therefore, although molecules can be represented by biological sequence structures, and previous models have also established evaluation indicators from the level of string similarity, this article does not involve these indicators in comparison, including NLP indicators such as BLEU, Levenshtein.
 Therefore, we use a series of metrics related to the similarity of molecular graphs:
@@
   • FTS: We employ three fingerprint metrics: MACCS FTS, RDK FTS, and Morgan FTS, where FTS stands for fingerprint Tanimoto similarity.
-MACCS (Durant et al.
-2002), RDK(Schneider, Sayle, and Landrum 2015) and Morgan(Rogers and Hahn 2010).
+MACCS [10], RDK[44] and Morgan[42].
 The fingerprints of two molecules are compared using Tanimoto similarity (also known as Jaccard index), and the average similarity over the evaluation dataset is reported.
 We use RDKIT toolkit.
-  • FCD score (Fréchet chemnet distance)(Preuer et al. 2018): Measure molecular similarity based on a pre-trained ”ChemNet” bioinformatics network. We use fcd 1.1 in python.
-D. Details for experiments
+  • FCD score (Fréchet chemnet distance): Measure molecular similarity based on a pre-trained "ChemNet" bioinformatics network. We use fcd 1.1 in python.
+For better evaluation, we also list some metrics which is not discussed in previous baseline. Here’re the information and results:
+  • Maximum Common Edge Substructure (MCES) : This metric finds the largest subgraph (in terms of the number of edges) that is common to two or more given graphs.
+We used the rdRascalMCES function in RDKit (Python) to measure MCES, with the default threshold set to 0.7.
+If either of the Johnson similarity estimates falls below this threshold, we consider as invalid for MCES.
+The results of MCES in CHEBI-20 is shown in Table 12.
+Method MCES
+BioT5 (reproduce) 0.8891
+gpt4-2024-04-09 (10 shot+BM25) 0.9060
+TGM-DLM 0.9036
+UTGDiff (our) 0.9166
+Table 12: MCES results in different models
+Appendix D Details for experiments
 Here’re the hyperparam for three different tasks describe in article, and here’s some general information on hyperparam:
@@
 We show the 3 seed example in forward reaction prediction in Table 17
-The specific data is listed in Table 11, 12, 13
+The specific data is listed in Table 14, 15, 16
 prompt
-Here we list the prompt for the three datasets: For CHEBI-20 dataset, We give its instruction form as : ”’[molecule description]’ is the description of molecule:” For the two editing dataset, since there’s task instruction in datasets, we don’t use any additional instructions.
-Here’s an example in Retrosynthesis dataset: ”Please suggest potential reactants for the given product.”; Here’s an example in forward reaction prediction dataset: ”With the provided reactants and reagents, propose a potential product.”
+Here we list the prompt for the three datasets: For CHEBI-20 dataset, We give its instruction form as : "’[molecule description]’ is the description of molecule:" For the two editing dataset, since there’s task instruction in datasets, we don’t use any additional instructions.
+Here’s an example in Retrosynthesis dataset: "Please suggest potential reactants for the given product."; Here’s an example in forward reaction prediction dataset: "With the provided reactants and reagents, propose a potential product."
+Also, here we list the prompt for large language model baselines. To limit the Generalist LLM to generate the desired SMILES, we use below prompts for both chatgpt and gpt-4 baseline:
+[…]
 Model MACCS $\uparrow$ RDK $\uparrow$ Morgan $\uparrow$ FCD $\downarrow$ Exact $\uparrow$ Valid $\uparrow$
@@
 Figure 6: The training curve in first 300K steps. Pretraining also demonstrates lower initial loss and faster convergence compared to training models from scratch.
-E. Source of datasets
+Appendix E Algorithm Pseudo Code
+The pseudo code of the training and sampling algorithm is shown in 1 and 2.
+Algorithm 1 Training Algorithm
+0: A graph $G=(V,E)$ with description $S$
+1:repeat
+2:  Sample $t\sim U(1,\cdot,T)$
+3:  Sample $G^{t}\sim V\bar{Q}_{V}^{t}\times E\bar{Q}_{E}^{t}$
+4:  $\hat{p}(S),\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
+5:  loss = $l_{CE}(\hat{p}(E),E)+l_{CE}(\hat{p}(V),V)+l_{CE}(\hat{p}(S),S)$
+6:  optimizer.step(loss)
+7:until converged
+Algorithm 2 Sampling Algorithm
+[…]
+2:for $t=T$ to 11 do
+3:  $\hat{p}(V),\hat{p}(E)=\phi_{\theta}(G^{t},S)$
+4:  $p_{\theta}(v^{t-1}_{i}\mid G^{t},S)=\sum_{v}q(v^{t-1}_{i}\mid v_{i}=v,v^{t}_{i})\;{\hat{p}}_{i}(v)$
+5:  $p_{\theta}(e^{t-1}_{ij}\mid G^{t},S)=\sum_{e}q(e^{t-1}_{ij}\mid e_{ij}=e,e^{t}_{ij})\;{\hat{p}}_{ij}(e)$
+6:  $G^{t-1}\sim\prod_{i}p_{\theta}(v^{t-1}_{i}\mid G^{t},S)\prod_{ij}p_{\theta}(e^{t-1}_{ij}\mid G^{t},S)$
+7:end for
+8:return  $G^{0}$
+Appendix F Potential Risks
+Potential risks associated with molecular generation models include the inadvertent creation of toxic or unstable compounds, which could pose significant safety hazards.
+Additionally, biases in the training data could be learned by the model, leading to unintended consequences.
+Another risk is the possibility of the model being used to discover harmful molecules instead of beneficial ones.
+These concerns highlight the necessity for stringent regulatory frameworks and ethical guidelines to govern the use of these models.
+Appendix G Source of codes and datasets
+Our Code is avaliable at: https://github.com/ran1812/UTGDiff
 All the datasets we used are open-sorcued, can be founded in github or huggingface: